### v1

# pQT_AMTs.bib

Abstract
Are there consistent and physically reasonable alternatives to the projection postulate? Does it have unique properties compared to acceptable alternatives? We answer these questions by systematically investigating hypothetical state-update rules for quantum systems that nature could have chosen over the Lüders rule. Among other basic properties, any prospective rule must define unique post-measurement states and not allow for superluminal signalling. Particular attention will be paid to consistently defining post-measurement states when performing local measurements in composite systems. Explicit examples of valid unconventional update rules are presented, each of them resulting in a distinct, well-defined foil of quantum theory. This framework of state-update rules allows us to identify operational properties that distinguish the projective update rule from all others and to put earlier derivations of the projection postulate into perspective.
1 Introduction
The main goal of this contribution is to identify operational properties that distinguish the quantum-mechanical projection postulate from other physically reasonable, but hypothetical, state-update rules. Justifications of Lüders rule have, not surprisingly, a long history, and a recent debate revolves around the claim that it can be derived from the other postulates of quantum theory (cf. Sec. 2.1). Many of these approaches have in common that their focus is on non-composite systems, either ignoring measurements on composite systems or making implicit assumptions about their consequences.
We study the nature of the projection postulate systematically by creating foil theories closely resembling quantum theory but equipped with different generalised state-update rules (GURs). In other words, we introduce generalised state-update theories (GUTs) for which the post-measurement states associated with a given outcome are not necessarily those prescribed by the Lüders rule. From the outset, we will take into account both single and composite systems. Alternative post-measurement rules have been considered occasionally (cf. Sec. 2.1), but we are not aware of any systematic approach.
To qualify as a legitimate update rule, the assignment of post-measurement states must satisfy a number of well-motivated physical requirements. The framework we present is broad enough to accommodate the standard quantum-mechanical projection postulate as well as many hypothetical rules, including the so-called passive measurements that leave the state of a system untouched while still producing outcome probabilities governed by Born’s rule.
In this setting, one can easily compare the properties of different update rules and, importantly, seek to identify those that are unique to the Lüders rule. It also provides a suitable backdrop to assess existing derivations of the Lüders rule summarised in Appendix A.
This paper is organised as follows. In Sec. 2 we set the scene by briefly reviewing earlier discussions related to the quantum mechanical state-update rule and describe the idea of quantum foil theories. Sec. 3 presents a simple axiomatic formulation of quantum theory and a detailed analysis of the state-update in composite quantum systems. In Sec. 4, we describe the basic requirements that state-update rules must satisfy and how they give rise to quantum foil theories. Explicit examples of legitimate and invalid update rules are given in Sec. 5. In Sec. 6 we explain what distinguishes the quantum-mechanical projection postulate from other valid update rules. In the concluding section, we summarise our results, discuss them, and outline future research.
2 Background and motivation
2.1 The quantum-mechanical state update rule
Using quantum systems as information carriers has led to applications without classical equivalents. Quantum measurements, in particular, provide a resource that contributes, for example, to secure communication protocols [bennett_quantum_2014], quantum teleportation [bennett_teleporting_1993] device-independent random number generators [colbeck_quantum_2009] and to speedy algorithms when using measurement-based quantum computation [briegel_measurement-based_2009].
The advantages of quantum measurements arise from fundamental differences when compared to measurements in classical theories. Classical measurements are deterministic in that they reveal pre-existing values of observables, and the effect on the state of the observed system is negligible, at least in principle. In quantum theory, measurement outcomes are probabilistic, and the state of a system after a measurement coincides with the pre-measurement state only in exceptional cases. Phenomenologically, the quantum mechanical state update is governed by the projection postulate, which describes a “collapse” of the state to an eigenstate of the operator representing the observed outcome. The collapse, also known as state reduction or Lüders rule [luders_uber_1950], ensures deterministic repeatability: two measurements of an observable always yield identical outcomes when implemented in quick succession [compton_directed_1925].
The standard postulates of quantum theory neither explain whether the state-update is a physical process [bohm_suggested_1952, dewitt_many-worlds_2015], nor do they define what constitutes a measurement [bell_against_1990, bacciagaluppi_heisenberg_2009]. The foundational questions resulting from this dilemma are often grouped under the term “measurement problem[busch_quantum_1996, norsen_foundations_2017]. The present work does not attempt to resolve the interpretational challenges posed by this long-standing problem.
Justifications of the Lüders rule have been given [bell_moral_1966, herbut_derivation_1969, hellmann_quantum_2016, kostecki_luders_2014, martinez_search_1990] (see Appendix A for a survey), and the difference between von Neumann’s and Lüders’ formulations of the projection postulate has been discussed repeatedly [vonneumann_mathematische_1932, luders_uber_1950, busch_luders_1998, sudbery_whose_2024]. Some authors have argued that the projection postulate follows from the other standard postulates of quantum theory [ozawa_quantum_1984, masanes_measurement_2019]; however, these derivations—one of which was the subject of a recent debate [stacey_masanes-galley-muller_2022, galley_reply_2022, kent_measurements_2023, masanes_response_2023]—appear to invariably rely on additional assumptions.
Occasionally, alternative post-measurement rules have been considered, either in the context of standard quantum theory [kent_nonlinearity_2005, kent_quantum_2021, masanes_measurement_2019] or of generalised probabilistic theories [kleinmann_sequences_2014]. A systematic investigation, as proposed here, has apparently not been carried out before.
2.2 Foil theories
Our approach is inspired by the constructive role of foil theories for the understanding of quantum theory. The strategy in defining foil theories is to embed quantum theory in a larger set of theories by relaxing specific assumptions about the structure of the state space, say, or the form of the time evolution. We briefly describe typical examples of foil theories and point out that a systematic investigation of generalised state-update rules has, apparently, not yet been carried out.
Generalised probabilistic theories (GPTs) [barrett_information_2007, hardy_quantum_2001, janotta_generalized_2014], for example, represent a large and versatile collection of foil theories. They are defined by their own sets of states, observables, and rules for system composition, all of which may differ from those of standard quantum theory. GPTs have become a standard tool to explore and identify distinctive characteristics of quantum theory which they contain as a special case. Many GPTs share properties such as the no-cloning theorem, uncertainty relations, or teleportation with quantum theory [barnum_cloning_2006, oppenheim_uncertainty_2010]. Some of them reproduce the standard quantum-mechanical correlations [wright_gleason-type_2019], while others exhibit superquantum correlations, e.g. Boxworld [popescu_quantum_1994].
Ontological models describe a conceptually different class of foil theories not necessarily based on the familiar Hilbert space structure of quantum theory. They aim to explain its predictions in terms of an observer’s incomplete knowledge about the ontic state that provides an objective and complete description of the state of a system. In these models, non-classical features of quantum theory such as complementarity of observables, superdense coding, interference phenomena, or uncertainty relations may result from epistemic restrictions [spekkens_evidence_2007, bartlett_reconstruction_2012].
Other foil theories differ from quantum theory in a single key feature only, such as attempts to modify Born’s rule [galvan_generalization_2008, galley_classification_2017]. The consequences of associating observables with non-Hermitian operators have been investigated, using PT-symmetric [bender_pt_2024] or normal [roberts_observables_2018] operators, for example. Quantum-like theories defined over the real numbers [stueckelberg_quantum_1960, renou_quantum_2021] or quaternions [finkelstein_foundations_1962] rather than the field of complex numbers have also been studied.
Nonlinear generalisations of Schrödinger’s equation have been suggested [beretta_quantum_1984, abrams_nonlinear_1998, ferrero_nonlinear_2004, kent_nonlinearity_2005, rembielinski_nonlinear_2020], most notably in Weinberg’s work [weinberg_testing_1989]. Gisin et al. [gisin_weinbergs_1990, simon_no-signaling_2001, bassi_no-faster-than-light-signaling_2015] pointed out that they imply violations of the no-signalling principle, although some specific nonlinear transformations were later found not to be ruled out by Gisin’s argument [czachor_nonlocal_1998, kent_nonlinearity_2005, ferrero_nonlinear_2004, helou_extensions_2017, rembielinski_nonlinear_2020].
In the specific context of modifications to the projection postulate, “causal quantum theory[kent_causal_2005] has been proposed, positing that the Lüders collapse is a well-defined physical process that satisfies strict local causality. Additionally, Kent introduced a new type of hypothetical measurement that is capable of revealing complete or partial information about the “local” state of a system, without actually inducing a state update [kent_nonlinearity_2005, kent_quantum_2021, kent_measurements_2023], an idea first presented in [busch_is_1997]. Such a measurement would, effectively, allow for yet other nonlinear state transformations without enabling superluminal signalling.
3 Quantum theory and the projection postulate
3.1 Axioms of quantum theory
The foil theories we introduce will differ from quantum theory only in their update rule assigning post-measurement states. The other standard postulates of non-relativistic, finite-dimensional quantum theory remain unchanged.
  • ($\mathsf{S}$)
    The states of a quantum system correspond to density operators $\rho\in\mathcal{S}(\mathcal{H})$, non-negative operators with unit trace acting on a finite-dimensional, separable and complex Hilbert space $\mathcal{H}$ associated with the system.
  • ($\mathsf{T}$)
    The time evolution of states, as well as any other reversible transformation, is described by a unitary operator $U\in\mathcal{U}(\mathcal{H})$.
  • ($\mathsf{C}$)
    The state space of a composite system is obtained from tensoring the state spaces describing its constituents, i.e. by $\mathcal{S}_{AB}=\mathcal{S}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$. Given a state $\rho_{AB}$ of the composite system, the state of subsystem $A$ is obtained by tracing out the other one, $\rho_{A}=\text{Tr}_{B}\left[\rho_{AB}\right]$, etc.
  • ($\mathsf{O}$)
    Observable quantities are represented by collections $M=\{P_{x}\}_{x=1}^{n}$ of mutually orthogonal projectors $P_{x}\in\mathcal{P}(\mathcal{H})$, $P_{x}P_{y}=P_{x}\delta_{xy}$, that sum to the identity operator on $\mathcal{H}$, i.e. $\sum_{x}P_{x}=\mathbb{I}$. Each projector $P_{x}$ represents a possible outcome of a measurement of the observable $M$.
  • ($\mathsf{B}$)
    The probability that a measurement of $M$ performed on a system in the state $\rho$ yields the outcome represented by $P_{x}$ is given by the Born rule: $\text{prob}(x)=\text{Tr}\left[P_{x}^{A}\,\rho\right]$.
Finally, the projection postulate describes what happens to the state of a quantum system when an observable is measured.
  • ($\mathsf{L}$)
    If a measurement of $M$ yields an outcome represented by the projection $P_{x}$, then the initial state $\rho\in\mathcal{S}(\mathcal{H})$ will update in the following way,
    $\rho\;\stackrel{{\scriptstyle x}}{{\longmapsto}}\;\rho_{x}=\frac{P_{x}\rho P_{x}}{\text{Tr}\left[P_{x}\rho P_{x}\right]}\in\mathcal{S}(\mathcal{H})\,.$ (1)
The focus of our work is to investigate the role of postulate ($\mathsf{L}$) by replacing it with alternative update rules. In view of later developments, it will be convenient to suppress the normalisation factor in Eq. (1). We introduce the map
$\rho\;\stackrel{{\scriptstyle x}}{{\longmapsto}}\omega^{\mathsf{L}}(P_{x},\rho)=P_{x}\rho P_{x}\,,$ (2)
which sends $\rho$ to a sub-normalized state [^1] A sub-normalised state is a non-negative operator on a Hilbert space with trace less than or equal to $1$, i.e. $\rho\in\bar{\mathcal{S}}(\mathcal{H})=\{\lambda\rho:\lambda\in[0,1],\rho\in\mathcal{S}(\mathcal{H})\}$. in the space $\bar{\mathcal{S}}(\mathcal{H})$, and call it the Lüders rule, giving postulate ($\mathsf{L}$) its name. The trace of the post-measurement state, $\text{Tr}(\omega^{\mathsf{L}}(P_{x},\rho))=\text{Tr}(P_{x}\rho)$, yields the outcome probability as prescribed by the Born rule of postulate ($\mathsf{B}$). In addition, the maps (1) and Eq. (2) do not require the input state to be normalised. Replacing $\rho$ by $\lambda\rho$, $\lambda\in[0,1]$, shows that the Lüders rule is 1-homogeneous: $\omega^{\mathsf{L}}(P_{x},\lambda\rho)=\lambda\omega^{\mathsf{L}}(P_{x},\rho)$.
The alternative update rules $\omega(P_{x},\rho)$ considered in Sec. 5 will be modeled on the Lüders rule in the sense that they are 1-homogeneous, only depend on the measurement outcome $P_{x}$ and the initial state $\rho$, and agree with the Born rule for outcome probabilities. Further physically motivated properties that any acceptable update rule should satisfy will be discussed in Sec. 4.
Before defining update rules replacing ($\mathsf{L}$), we must take a closer look at the effect of quantum measurements performed on composite systems. As it stands, the rule ($\mathsf{L}$) does, in fact, not specify the post-measurement state of a composite system when measurements are performed on subsystems. The post-measurement state traditionally attributed to a quantum system relies on an additional implicit assumption.
3.2 Measurements on composite quantum systems
The projection postulate ($\mathsf{L}$) does not explicitly cover local measurements which arise in composite quantum systems, i.e. measurements that are carried out on subsystems. To discuss this situation in detail, we will consider a system with two constituents $A$ and $B$, say, and an associated Hilbert space $\mathcal{H}_{AB}=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, in line with Axiom ($\mathsf{C}$).
Suppose we wish to measure the global observable $M_{AB}$, represented by an operator acting on the product space $\mathcal{H}_{AB}$. To this effect, a global device $\mathcal{D}_{AB}$ must be used that interacts with the composite system in its entirety. The outcome of the measurement is represented by one of the projection operators $P_{x}^{AB}$ (associated with $M_{AB}$), that sum to the identity on the space $\mathcal{H}_{AB}$, as described in Axiom ($\mathsf{O}$). The probability to obtain any of these outcomes is found from Axiom ($\mathsf{B}$), and the state $\rho_{AB}$ of the overall system will update according to the Lüders rule ($\mathsf{L}$) given in Eq. (1). The projectors $P_{x}^{AB}$ do not have to respect the product structure of the space $\mathcal{H}_{AB}$: for a two-qubit system, they may project onto the states of the Bell basis, for example. At no point do global measurements refer to the product structure of the space $\mathcal{H}_{AB}$.
However, on a bipartite quantum system, one may also carry out local measurements. They arise naturally when its constituents are in separate locations. Local measurements are implemented by measuring devices $\mathcal{D}_{A}$ or $\mathcal{D}_{B}$ that act on subsystems only, i.e. the constituents $A$ or $B$, respectively. The postulate ($\mathsf{L}$) does not cover this situation: no post-measurement state $\rho_{AB,x}\in\mathcal{S}_{AB}\equiv\mathcal{S}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ is specified by ($\mathsf{L}$) when using the device $\mathcal{D}_{A}$ to carry out a measurement of the local observable $M_{A}$, resulting in the outcome $x$. Let us explain why.
According to axiom ($\mathsf{O}$), the local observable $M_{A}$ corresponds to a collection of mutually orthogonal projectors $\{P_{x}^{A}\}_{x=1}^{n}$, defined on $\mathcal{H}_{A}$ and summing to the identity, $\sum_{x}P_{x}^{A}=\mathbb{I}_{A}$. Thus, we associate measurement outcomes of $M_{A}$ with projectors $P_{x}^{A}$. However, neither Eq. (1) nor (2) can be used to determine the post-measurement state of the system as a whole: to apply Lüders rule, we need to use (global) projectors $P_{x}^{AB}$ defined on $\mathcal{H}_{AB}$ but the (local) operators $P_{x}^{A}$ act on $\mathcal{H}_{A}$ only.
For the experimenter carrying out a local measurement on subsystem $A$, the existence of the other constituent $B$ is irrelevant. If the composite system resides in the state $\rho_{AB}$, we invoke Axiom ($\mathsf{C}$): the initial state of subsystem $A$ is given by the density matrix $\rho_{A}$ obtained by tracing out subsystem $B$. Then, we are in a position to apply ($\mathsf{L}$) to find the post-measurement state of subsystem $A$ upon obtaining the local outcome $P_{x}^{A}$, namely
$\rho_{A}\;\stackrel{{\scriptstyle x}}{{\longmapsto}}\rho_{A,x}\propto\;\omega^{\mathsf{L}}(P_{x}^{A},\rho_{A})\,,$ (3)
where the operators $\rho_{A}=\text{Tr}_{B}(\rho_{AB})$ and $P_{x}^{A}$ are defined on the same Hilbert space $\mathcal{H}_{A}$, as required by the Lüders rule. The relation (3) does, of course, not mean that the system as a whole now resides in a product state. Importantly, when obtaining a specific outcome in a single run, Eq. (3) says nothing about the post-measurement state of the other subsystem, or of the system as a whole. [^2] It is only possible to conclude that the pre-measurement state $\rho_{AB}$ assigns a non-zero probability to obtaining the outcome $P^{A}_{x}$. Read in this way, the statement of Lüders rule given by ($\mathsf{L}$) is incomplete as it does not fully prescribe how to update the state of a quantum system when a local measurement is performed. [^3] The update rule also seems to be incomplete if one were to use a different formulation of ($\mathsf{L}$) stating that, after a measurement, the system resides in an eigenstate of the measured observable.
There are, however, consistency requirements when performing measurements on a composite system. If the outcome $P_{x}^{A}$ is non-degenerate, then the Lüders rule assigns a pure post-measurement state $\omega^{\mathsf{L}}(P_{x}^{A},\rho_{A})$ to subsystem $A$. Consequently, the post-measurement state of the composite system must be a product state, whatever the initial state $\rho_{AB}$. On its own, the projection postulate ($\mathsf{L}$) provides, however, no information about the local state at $B$ upon performing a local measurement of $M_{A}$ on the constituent $A$. In principle, any density matrix $\rho_{B}=\text{Tr}_{B}(\rho_{AB})\in\mathcal{S}(\mathcal{H}_{B})$ is compatible. For example, one could imagine that any local measurement at $A$ throws the state of the composite system into a product state, with the post-measurement state at $B$ being equal to the maximally mixed state $\rho_{B}\propto\mathbb{I}_{B}$.
The correct global post-measurement state after a local measurement on a quantum system is, of course, well-known. Calculating quantum correlations in Bell-type experiments relies on knowing the state of subsystem $B$ if the outcome of a local measurement at $A$ has been found. [^4] Recall that, mathematically, the state that correctly reproduces the experimental findings is obtained in the following way. Write the initial state corresponding to $\rho_{AB}$ as a superposition of product terms, using the eigenstates of the operator $M_{A}$ for the first factor. Then, identify the term with the label $x$, the value of the observed measurement outcome in a given run. The second factor in this expression characterises the correct post-measurement state for subsystem $B$. Formally, upon measuring the local observable $M_{A}$ and finding $P_{x}^{A}$ given $\rho_{AB}\in\mathcal{S}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$, the state of the composite system updates in this way,
$\rho_{AB}\;\stackrel{{\scriptstyle x}}{{\longmapsto}}\;\rho_{AB,x}=\frac{P_{x}^{A}\otimes\mathbb{I}_{B}\,\rho_{AB}\,P_{x}^{A}\otimes\mathbb{I}_{B}}{\text{Tr}\left[P_{x}^{A}\otimes\mathbb{I}_{B}\,\rho_{AB}\,P_{x}^{A}\otimes\mathbb{I}_{B}\right]}\in\mathcal{S}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})\,.$ (4)
This rule generalises the postulate ($\mathsf{L}$) to cover the update of the global state $\rho_{AB}$ caused by a local measurement,
$\rho_{AB}\;\stackrel{{\scriptstyle x}}{{\longmapsto}}\;\omega^{\mathsf{L}}(P_{x}^{A},\rho_{AB})=P_{x}^{A}\otimes\mathbb{I}_{B}\,\rho_{AB}\,P_{x}^{A}\otimes\mathbb{I}_{B}\in\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})\,.$ (5)
Stated differently, we are making the point that the projection postulate ($\mathsf{L}$) or, equivalently, Lüders rule (2), does not say how the state of a composite system updates under local measurements. Additional experimental input is required to arrive at the rule (4). For simplicity, we continue to use the symbol $\omega^{\mathsf{L}}$ to denote the complete update rule which defines the effects of both local and global measurements.
The update rule (4) does have a simple physical interpretation. Suppose we use a device $\mathcal{D}_{AB}$ to perform a global measurement of the observable $M_{A}\otimes\mathbb{I}_{B}$, with outcomes represented by the projector $P_{x}^{A}\otimes\mathbb{I}_{B}$. To find the post-measurement state of the composite system, we can apply the original form of the projection postulate ($\mathsf{L}$), only to find the update rule (4).
The distinction between a local measurement of $M_{A}$ implemented by the device $\mathcal{D}_{A}$ and a global one of $M_{A}\otimes\mathbb{I}_{B}$, effected by the device $\mathcal{D}_{AB}$, will be important to express the requirements that alternatives to the Lüders rule must satisfy.
4 Generalising the state-update rule
4.1 Preliminaries
In this section, we will define the properties that an acceptable state-update rule must have. The main role of an update rule is, of course, to determine the post-measurement state of the measured system. We need to allow that measurements on a part of a composite system may affect the state of other subsystems not acted upon. This behaviour is known to occur in quantum theory and gave rise to the discussion in Sec. 3.2. In other words, we must ensure that an update rule assigns post-measurement states consistently to composite systems when local measurements are performed. For convenience, we will focus on bipartite systems; the extension to multipartite systems is straightforward. [^5] Anticipating the discussion of Sec. 4.3 on the role proper and improper mixed states, we assume for now that the density operators we consider represent either pure states or improper mixed states, i.e. those arising from entanglement with other systems.
A generic update rule will be denoted by the letter $\omega$, with a subscript identifying the possibly composite system to which it applies. Let us consider a composite system with Hilbert space $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, initially residing in the state $\rho_{AB}\in\mathcal{S}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$. If a local measurement is performed on subsystem $\mathcal{H}_{A}$ and yields the outcome $P_{x}^{A}$, the operator $\omega_{AB}(P_{x}^{A},\rho_{AB})$ will represent the sub-normalised post-measurement state of the system, in analogy with the Lüders rule of Eq. (5). [^6] A different formula is generally required to describe the update of proper mixed states, i.e. those due to the incomplete knowledge about the preparation of a system (cf. Sec. 4.3). The subsystem on which the measurement is performed can be deduced from the Hilbert space associated with the projector $P_{x}^{A}$ characterising the measurement outcome. In the current example, it is the system with label $A$.
Any update rule must conform with postulate ($\mathsf{B}$) implying a non-trivial constraint. By taking the trace of a sub-normalised post-measurement state, the Born rule prescribes the probability to obtain outcome $P_{x}^{A}$,
$\text{Tr}\left[\omega_{AB}\left(P_{x}^{A},\rho_{AB}\right)\right]=\text{Tr}\left(P_{x}^{A}\otimes\mathbb{I}_{B}\,\rho_{AB}\right)=\text{Tr}\left(P_{x}^{A}\,\text{Tr}_{B}\left(\rho_{AB}\right)\right)\,.$ (6)
An analogous condition is, of course, assumed for measurements on subsystem $\mathcal{H}_{B}$.
When considering sequences of measurements, 1-homogeneity, described following Eq. (2) for the quantum-mechanical update rule, is a convenient mathematical property. It means that an update rule satisfies the constraint
$\omega_{AB}\left(P_{x}^{A},\lambda\,\rho_{AB}\right)=\lambda\,\omega_{AB}\left(P_{x}^{A},\rho_{AB}\right)\,,\quad\lambda\in\left[0,1\right]\,,$ (7)
for all initial states $\rho_{AB}\in\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ and measurement outcomes $P_{x}^{A}\in\mathcal{P}(\mathcal{H}_{A})$. This property allows us to concatenate update rules without the need to repeatedly normalise post-measurement states. Now suppose that two time-ordered measurements with outcomes $P_{x}$ and $P_{y}$ are performed on a composite system in state $\rho_{AB}$. Then, the final state is simply given by $\omega_{AB}\left(P_{y},\omega_{AB}\left(P_{x},\rho_{AB}\right)\right)$. Here the outcomes $P_{x}$ and $P_{y}$ may correspond to measurements performed on the same subsystem or on different subsystems.
4.2 Necessary properties of generalised update rules
We will now list physically or operationally motivated requirements for update rules. These assumptions will lead to a concise definition of a generalised update rule, given in Def. 1. Naturally, these requirements are satisfied by the quantum mechanical Lüders rule.
The requirements will take into account that update rules must hold for single and composite systems. For simplicity, we will usually consider only measurements on one subsystem, say $\mathcal{H}_{A}$, though analogous conditions are assumed to apply to measurements on any subsystem.
The first three assumptions do not explicitly rely on the composite structure. Completeness requires that every conceivable measurement outcome specifies a unique post-measurement state. Context-independence asserts that the observed outcome and the pre-measurement state alone determine the post-measurement state. It introduces a form of non-contextuality [kochen_problem_1967] that is present in quantum theory—namely, that all operationally equivalent experimental runs (i.e. involving the same initial state and the same measurement outcome) yield identical post-measurement states. Mathematically, completeness and context-independence determine the co-domain and the domain of the update rule $\omega_{AB}$ for a composite system, respectively.
(A1-A2)
Completeness &\&& context-independence A complete and context-independent update rule maps any combination of a (local) outcome, $P_{x}^{A}\in\mathcal{P}(\mathcal{H}_{A})$, and a sub-normalised state, $\rho_{AB}\in\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$, to another sub-normalised state,
$\omega_{AB}:\mathcal{P}(\mathcal{H}_{A})\times\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})\to\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})\,.$ (8)
Thirdly, local covariance ensures that different observers provide consistent descriptions of the same experiment. For a measurement on a non-composite system, suppose Alice’s description involves the state $\rho$ and outcome $P_{x}$. In Bob’s reference frame, the same measurement will be described by the state $U\rho U^{\dagger}$ and the outcome $UP_{x}U^{\dagger}$, where $U$ represents the frame change. Local covariance ensures that Alice’s and Bob’s post-measurement states remain related by $U$, preventing any contradictory predictions concerning future measurements. Effectively, we require that, instead of applying a unitary $U$ on a system after measuring an observable $\{P_{x}\}_{x}$, one can equivalently implement it before measuring the suitably transformed observable $\{UP_{x}U^{\dagger}\}_{x}$. The term “local” gains particular significance when considering the more general scenario of a measurement performed on a subsystem of a composite system. To ensure consistency across descriptions of a local measurement, the same joint state must be obtained when applying the unitary transformation $U_{A}\otimes U_{B}$ (i) after measuring $\mathcal{H}_{A}$ with outcome $P_{x}^{A}$, or (ii) before measuring $\mathcal{H}_{A}$ with rotated outcome $U_{A}P_{x}^{A}U_{A}^{\dagger}$. Local unitaries $U_{A}\otimes U_{B}$ represent changes to frames that preserve the composite structure of $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$.
Local covariance For all measurement outcomes $P_{x}^{A}\in\mathcal{P}(\mathcal{H}_{A})$, states $\rho_{AB}\in\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ of the joint system, and local transformations $U_{A}\in\mathcal{U}(\mathcal{H}_{A})$ and $U_{B}\in\mathcal{U}(\mathcal{H}_{B})$, we must have
$\omega_{AB}\left(U_{A}P_{x}^{A}U_{A}^{\dagger},U_{A}\otimes U_{B}\rho_{AB}U_{A}^{\dagger}\otimes U_{B}^{\dagger}\right)=\left(U_{A}\otimes U_{B}\right)\left(\omega_{AB}\left(P_{x}^{A},\rho_{AB}\right)\right)\left(U_{A}^{\dagger}\otimes U_{B}^{\dagger}\right)\,.$ (9)
The remaining assumptions rely on the composite structure of the system. We require self-consistency of the update rule: the assigned post-measurement states must not depend on whether or not the probed system is regarded as part of a larger composite system. This property ensures that the assignment of post-measurement states for composite systems is unambiguous. Violating self-consistency would result in an inconsistent treatment of measurements in the framework, wherein an observer’s choice of whether to regard a system as a subsystem (and how to partition its environment) would have measurable effects. Mathematically, self-consistency is ensured by the requirement that the state of any system after the measurement could also be obtained by applying $\omega$ to a larger system and tracing out the irrelevant subsystems.
Self-consistency Suppose that the system with label $B$ is, in fact, a composite system, i.e. $\mathcal{H}_{B}=\mathcal{H}_{B^{\prime}}\otimes\mathcal{H}_{B^{\prime\prime}}$, hence the update rule can be relabeled as $\omega_{AB}=\omega_{AB^{\prime}B^{\prime\prime}}$. For all $\mathcal{H}_{B^{\prime\prime}}$, joint states $\rho_{AB^{\prime}B^{\prime\prime}}\in\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B^{\prime}}\otimes\mathcal{H}_{B^{\prime\prime}})$ and measurement outcomes $P_{x}\in\mathcal{P}(\mathcal{H}_{A})$, we must have
$\text{Tr}_{B^{\prime\prime}}\left[\omega_{AB^{\prime}B^{\prime\prime}}\left(P_{x}^{A},\rho_{AB^{\prime}B^{\prime\prime}}\right)\right]=\omega_{AB^{\prime}}\left(P_{x}^{A},\text{Tr}_{B^{\prime\prime}}\left(\rho_{AB^{\prime}B^{\prime\prime}}\right)\right)\,.$ (10)
Self-consistency entails that the update rule $\omega_{A}$ of a non-composite system emerges as a special case of the composite-system update rule $\omega_{AB}$: letting $\rho_{A}=\text{Tr}_{B}(\rho_{AB})$, we have
$\omega_{A}\left(P_{x}^{A},\rho_{A}\right)=\text{Tr}_{B}\left(\omega_{AB}\left(P_{x}^{A},\rho_{AB}\right)\right)\,.$ (11)
Any update rule must respect the (quantum) no-signalling principle, which prohibits space-like separated parties from communicating through local measurements. Mathematically, we require that unconditional measurements—where no specific outcome is selected—do not alter the reduced state of unmeasured subsystems.
No-signalling For all joint states $\rho_{AB}\in\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ and all local observables $\{P_{x}^{A}\}_{x}$ on $\mathcal{H}_{A}$, we must have
$\sum_{x}\text{Tr}_{A}\left[\omega_{AB}\left(P_{x}^{A},\rho_{AB}\right)\right]=\text{Tr}_{A}\left(\rho_{AB}\right)\,.$ (12)
Finally, we assume local commutativity: conditioned on fixed outcomes for measurements on different subsystems, the order in which these measurements are performed does not matter. As with no-signalling, local commutativity finds justification in special relativity. Suppose that local measurements on two space-like separated subsystems yield outcomes $P_{x}^{A}$ and $P_{y}^{B}$. Since the time-ordering of the two measurements is relative, we require that the same post-measurement state of the combined system is obtained irrespective of whether $P_{x}^{A}$ or $P_{y}^{B}$ is measured first. This property also implies that the post-measurement state of the combined system is defined unambiguously if $P_{x}^{A}$ and $P_{y}^{B}$ are measured simultaneously.
Local commutativity For all measurement outcomes $P_{x}^{A}\in\mathcal{P}(\mathcal{H}_{A})$, $P_{y}^{B}\in\mathcal{P}(\mathcal{H}_{B})$ and joint states $\rho_{AB}\in\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$, we must have
$\omega_{AB}\left(P_{x}^{A},\omega_{AB}\left(P_{y}^{B},\rho_{AB}\right)\right)=\omega_{AB}\left(P_{y}^{B},\omega_{AB}\left(P_{x}^{A},\rho_{AB}\right)\right)\,.$ (13)
Requirements A1-A6 are the core assumptions feeding into the framework of generalised state-update theories. Other features of the Lüders rule of quantum theory will not necessarily be shared by generalised update rules. For example, the resulting foil theories do not have to be deterministically repeatable, and they do not have to allow for quantum measurements to preserve the indistinguishability of preparations (cf. Sec. 4.3). As we will see in Sec. 5, there are generalised state-update theories (GUTs) in which different preparations of the same mixed state do not remain indistinguishable after a measurement. Thus, depending on the update rule, density operators may or may not provide a complete description of an individual system.
We are now in a position to present the definition of generalised state-update rules.
Definition 1.
A generalised state-update rule $\omega$ is a set of functions $\{\omega_{AB}\}$—where each function $\omega_{AB}$ is associated with a pair of Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$—satisfying the following properties: (i) consistency with the Born rule (Eq. (6)); (ii) 1-homogeneity (Eq. (7)); and (iii) assumptions A1-A6 (Eqs. (8)-(13)).
Any update rule chosen to replace the projection postulate of quantum theory will give rise to one specific generalised state-update theory. Multiple update rules may induce the same behaviour on non-composite systems via Eq. (11). This point is illustrated in Sec. 5.2 where different extensions of the single-system Lüders rule to multi-partite systems are presented.
How do the update rules defined here compare with quantum instruments [heinosaari_mathematical_2011], i.e. collections $\{\mathcal{I}_{x}\}_{x}$ of quantum operations [^7] Quantum operations are linear, completely positive and trace non-increasing maps defined on the space $\mathcal{L}(\mathcal{H})$ of bounded operators acting on $\mathcal{H}$. Quantum instruments are usually defined as mappings from an outcome space $\left(X,\Sigma\right)$ to the set of quantum operations. However, since we will only deal with discrete observables, an instrument is completely determined by the finite set of operations $\{\mathcal{I}_{x}\}_{x}$. that describe the effect on states of quantum measurements? The main difference is that the state-update rules satisfying our definition are not required to be linear or completely positive. Consequently, the set of alternative state-update rules is larger than the set of quantum instruments.
4.3 Updating (im-)proper mixtures
For a pure joint state $\rho_{AB}$, Eq. (11) applies irrespective of $\rho_{AB}$ being a product state or being entangled. Accordingly, $\omega_{A}\left(P_{x}^{A},\rho_{A}\right)$ represents the post-measurement state of system $\mathcal{H}_{A}$ when $\rho_{A}$ is either a pure state or an improper mixture, i.e. the reduced density operator of an entangled state. In contrast, proper mixtures result from classical uncertainty about the “true” preparation of a system. A proper mixture can be thought of as a Gemenge (German for ‘mixture’), i.e. as a collection $\mathcal{G}=\{(p_{1},\rho_{1}),(p_{2},\rho_{2}),\dots,(p_{n},\rho_{n})\}$ where $p_{i}\geq 0$, $\sum_{i}p_{i}=1$ and $\rho_{i}\in\bar{\mathcal{S}}(\mathcal{H})$ represent pure states or improper mixed states. The Gemenge $\mathcal{G}$ describes a system that is prepared in state $\rho_{i}$ with probability $p_{i}$ [busch_quantum_1996]. To any Gemenge there corresponds a unique density operator, $\rho_{\mathcal{G}}=\sum_{i=1}^{n}p_{i}\rho_{i}$.
If a system described by $\mathcal{G}$ is measured with outcome $P_{x}^{A}$, the updated description is given by the Gemenge $\mathcal{G}_{x}$ with the same weights $p_{i}$ but with updated (sub-normalised) states, $\mathcal{G}_{x}=\{p_{i},\omega_{A}(P_{x}^{A},\rho_{i})\}$. In other words, if the system resided in the state $\rho_{i}$ with probability $p_{i}$ before the measurement with outcome $P_{x}^{A}$, it will reside in the state $\omega_{A}(P_{x}^{A},\rho_{i})$ with probability $p_{i}$ after the measurement. [^8] If the outcome $P_{x}^{A}$ is never observed, $\text{Tr}(P_{x}^{A}\rho_{i})=0$, then $\omega_{A}(P_{x}^{A},\rho_{i})=O$, where $O$ is the zero operator, and the experimenter learns that the system was not prepared in the state $\rho_{i}$. Thus, the density matrix updates according to the rule
$\rho_{\mathcal{G}}=\sum_{i=1}^{n}p_{i}\rho_{i}\quad\stackrel{{\scriptstyle P_{x}^{A}}}{{\mapstochar\relbar\joinrel\relbar\joinrel\relbar\joinrel\rightarrow}}\quad\rho_{\mathcal{G}_{x}}=\sum_{i=1}^{n}p_{i}\;\omega_{A}\left(P_{x}^{A},\rho_{i}\right)\,.$ (14)
A similar expression applies for local measurements on a subsystem when the composite system is described by a proper mixture, i.e. $\mathcal{G}=\{(p_{1},\rho^{1}_{AB}),(p_{2},\rho^{2}_{AB}),\dots,(p_{n},\rho^{n}_{AB})\}$.
The update rule is therefore applied differently depending on whether a mixed state is regarded as a proper or an improper mixture. As shown in (14), in the case of a proper mixture, the map acts on each operator $\rho_{i}$ appearing in the decomposition of $\rho_{\mathcal{G}}$ and included in the associated Gemenge $\mathcal{G}$. Typically, the post-measurement state $\rho_{\mathcal{G}_{x}}$ also represents a proper mixture. Consequently, sequences of measurements on a proper mixture are described by repeatedly applying the update rule to each pure or improper mixed state $\rho_{i}$ in the initial Gemenge $\mathcal{G}$.
The equations (8)-(13) formalising assumptions A1-6 were formulated with pure and improper mixed states in mind. Nevertheless, they also ensure a consistent treatment of measurements on proper mixtures. For instance, the completeness A1 of $\omega$ guarantees that $\rho_{\mathcal{G}_{x}}$ in (14) is a valid state of the system. Similar arguments apply to assumptions A3-6, with the exception of context-independence A2, which requires minor adaptation. In fact, the post-measurement state $\rho_{\mathcal{G}_{x}}$ generally does not depend on the initial mixed state $\rho_{\mathcal{G}}$, but rather on the initial Gemenge $\mathcal{G}$. When proper mixtures are included, context-independence is generalised as follows: the observed outcome and the pre-measurement Gemenge alone determine the post-measurement state. [^9] If $\rho_{A}$ denotes either a pure or an improper mixed state, we can associate the trivial Gemenge $\mathcal{G}=\{(1,\rho_{A})\}$, describing a system prepared in the state $\rho_{A}$ with probability $1$.
By considering proper and improper mixtures separately, we allow the framework to include foil theories where sequential measurements can be used to distinguish between different realisations of the same mixed state. Unlike quantum theory, such theories do not generally preserve the indistinguishability of preparations. In addition, our framework should also include update rules that send pure states to proper mixtures, thus introducing another level of uncertainty into the behaviour of a measuring device. In such cases, equations such as (13) for local commutativity A6, which involve sequential measurements, must be adjusted by ensuring that the second update rule is applied to each state in the Gemenge resulting from the first measurement.
4.4 Foil theories from generalised state-update rules
Having defined update rules in Def. 1, we now introduce GUTs.
Definition 2.
A Generalised state-Update Theory (GUT) is defined by the quantum postulates for states ($\mathsf{S}$), time evolution ($\mathsf{T}$), system composition ($\mathsf{C}$), observables ($\mathsf{O}$), outcome probabilities ($\mathsf{B}$), and by a generalised state-update rule $\omega$ assigning post-measurement states.
5 Examples of generalised state-update rules
We now introduce several examples of update rules and investigate properties they have in addition to the essential requirements A1-A6 of Sec. 4.2. Each of the hypothetical rules defines a foil of quantum theory through Def. 2 in Sec. 4.4, characterised by a modified projection postulate.
5.1 Lüders measurements: quantum theory
Our framework includes quantum theory, since the properties A1-A6 underlying Def. 1 for generalised update rules have been abstracted from the Lüders rule,
$\omega^{\mathsf{L}}_{AB}(P_{x}^{A},\rho_{AB})=P_{x}^{A}\otimes\mathbb{I}_{B}\rho_{AB}P_{x}^{A}\otimes\mathbb{I}_{B}\,.$ (15)
In Sec. 3.2, we described an important property of the Lüders rule, that we will call composition compatibility. A local measurement on subsystem $\mathcal{H}_{A}$, implemented by a device $\mathcal{D}_{A}$ and yielding the outcome $P_{x}^{A}$, leads to the same post-measurement state of the composite system as a global measurement on $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, implemented by a device $\mathcal{D}_{AB}$ and yielding the outcome represented by $P_{x}^{A}\otimes\mathbb{I}_{B}$.
Definition 3 (Composition Compatibility).
An update rule $\omega$ satisfies composition compatibility if, for every $\mathcal{H}_{A}$, $\mathcal{H}_{B}$, joint state $\rho_{AB}\in\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ and local outcome $P_{x}^{A}\in\mathcal{P}(\mathcal{H}_{A})$,
$\omega_{AB}\left(P_{x}^{A},\rho_{AB}\right)=\omega_{AB}\left(P_{x}^{A}\otimes\mathbb{I}_{B},\rho_{AB}\right)\,.$ (16)
We will see that not all update rules satisfy this property—in other words, it does not follow from assumptions A1-A6. Note that the two expressions in Eq. (16) capture two different situations: the function on the left-hand side is defined on $\mathcal{P}(\mathcal{H}_{A})\times\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$, while that on the right-hand side on $\mathcal{P}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})\times\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$. In physical terms, one side of Eq. (16) corresponds to a local measurement on $\mathcal{H}_{A}$, while the other side is associated with a global measurement on $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. Eq. (16) refers to measurements on $\mathcal{H}_{A}$ only, but the property is assumed to hold for measurements on any subsystem.
Composition compatibility (16) can therefore be regarded as the additional assumption that extends the “operationally incomplete” postulate ($\mathsf{L}$) for single systems, i.e.
$\omega^{\mathsf{L}}_{A}(P_{x}^{A},\rho_{A})=P_{x}^{A}\rho_{A}P_{x}^{A}\,,$ (17)
to the Lüders rule $\omega^{\mathsf{L}}_{AB}$ for multi-partite systems, Eq. (15), which encompasses the effects of both local and global measurements.
Another argument to extend the single-system Lüders rule to composite systems uses the fact that $\omega^{\mathsf{L}}_{A,x}(\cdot)\equiv\omega^{\mathsf{L}}_{A}(P^{A}_{x},\cdot):\bar{\mathcal{S}}(\mathcal{H}_{A})\to\bar{\mathcal{S}}(\mathcal{H}_{A})$, which denotes the Lüders rule for system $\mathcal{H}_{A}$ conditioned on the measurement outcome $P_{x}^{A}$, is convex-linear over $\bar{\mathcal{S}}(\mathcal{H}_{A})$. Therefore, $\omega^{\mathsf{L}}_{A,x}$ has a unique linear extension to the set of bounded operators $\mathcal{L}(\mathcal{H}_{A})$ (cf. [heinosaari_mathematical_2011]). The Lüders rule (15) for the composite system $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, conditioned on the local measurement outcome $P_{x}^{A}$, is then obtained by setting
$\omega^{\mathsf{L}}_{AB}\left(P_{x}^{A},\rho_{AB}\right)=\left(\omega^{\mathsf{L}}_{A,x}\otimes\mathcal{I}_{B}\right)\left(\rho_{AB}\right)\,$ (18)
for all $\rho_{AB}\in\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$, where $\left(\omega^{\mathsf{L}}_{A,x}\otimes\mathcal{I}_{B}\right)\left(L_{A}\otimes L_{B}\right)=\omega^{\mathsf{L}}_{A,x}\left(L_{A}\right)\otimes L_{B}$ for all $L_{A}\in\mathcal{L}(\mathcal{H}_{A})$ and $L_{B}\in\mathcal{L}(\mathcal{H}_{B})$. Here, $\mathcal{I}_{B}$ denotes the identity channel on $\mathcal{H}_{B}$. Eq. (18) is well-defined, as well as linear over $\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$, because $\omega^{\mathsf{L}}_{A,x}$ can be linearly extended to $\mathcal{L}(\mathcal{H}_{A})$. In particular, the completeness A1 of $\omega^{\mathsf{L}}_{AB}\left(P_{x}^{A},\rho_{AB}\right)$ follows from the complete positivity of $\omega^{\mathsf{L}}_{A,x}$: if $\omega^{\mathsf{L}}_{A,x}$ was not completely positive, then the operator $\omega^{\mathsf{L}}_{AB}\left(P_{x}^{A},\rho_{AB}\right)$ would not be positive for some $\rho_{AB}$.
Typically, generalised update rules have to neither satisfy deterministic repeatability nor composition compatibility. They may also lack some other properties such as preparation indistinguishability, which ensures that different preparations of the same mixed state remain indistinguishable after measurements. This property is equivalent to the statement that “density operators provide complete descriptions of quantum systems”. Moreover, ideality (measurement leaves a state unchanged if the outcome is certain) and local tomography (local measurements of $M_{A}$ and $M_{B}$ on different subsystems are operationally equivalent to a global measurement of the product observable $M_{A}\otimes M_{B}$ if the experimenters are allowed to communicate classically) may or may not be present. The Lüders rule is also endowed with a form of coherence, according to which the outcome probability distribution of an observable $M$ is not affected by a prior coarse-grained measurement of $M$ (see Def. 4 in Sec. 6). Table 1 summarises the extent to which these and other properties hold for the hypothetical update rules introduced in the following sections.
$\omega^{\mathsf{L}}$ $\omega^{\mathsf{locL}}$ $\omega^{\mathsf{P}}$ $\omega^{\mathsf{dep}}$ $\omega^{\lambda}$
deterministic repeatability $\times$ $\times$ $\times$
preparation indistinguishability (✓) $\times$ $\times$
composition compatibility $\times$ $\times$ $\times$
ideality (✓) $\times$ (✓)
local tomography $\times$ $\times$ $\times$
coherence $\times$ $\times$ $\times$
nonlocality $\times$ $\times$ $\times$
complete positivity $\times$ $\times$ $\times$
weak repeatability $\times$ $\times$
Table 1: Summary of the operational properties of the update rules discussed in Sec. 5. Ticks in parentheses indicate that the property holds for measurements on non-composite systems only.
5.2 Locally-Lüders measurements: correlation-free quantum theory
As noted in Sec. 4, the Lüders rule for single systems does not fix the state-update rule in multi-partite systems. Consequently, there will exist other update rules for composite systems that reduce to the single-system Lüders rule. One example is the 1-homogeneous map
$\omega_{AB}^{\mathsf{locL}}\left(P_{x}^{A},\rho_{AB}\right)=\frac{\omega^{\mathsf{L}}_{A}\left(P_{x}^{A},\text{Tr}_{B}\left(\rho_{AB}\right)\right)\otimes\text{Tr}_{A}\left(\rho_{AB}\right)}{\text{Tr}\left(\rho_{AB}\right)}$ (19)
which satisfies assumptions A1-A6 and thus defines a valid update rule. While manifestly different from the standard Lüders rule of Eq. (15), it assigns the same post-measurement states as postulate ($\mathsf{L}$) to non-composite systems. Using Eq. (11) and letting $\rho_{A}=\text{Tr}_{B}(\rho_{AB})$, we obtain
$\omega_{A}^{\mathsf{locL}}\left(P_{x}^{A},\rho_{A}\right)=\text{Tr}_{B}\left(\omega_{AB}^{\mathsf{locL}}\left(P_{x}^{A},\rho_{AB}\right)\right)=\omega_{A}^{\mathsf{L}}\left(P_{x}^{A},\rho_{A}\right)\,.$ (20)
In other words, $\omega^{\mathsf{locL}}$ represents an alternative extension of postulate ($\mathsf{L}$) but, unlike the Lüders rule $\omega^{\mathsf{L}}$, it does not satisfy composition compatibility, and is not convex-linear over pre-measurement states of the composite system.
The GUT defined by the update rule (19) is therefore locally indistinguishable from quantum mechanics but creates no correlations between entangled systems. Regardless of the observed outcome, a local measurement on subsystem $A$ leaves the reduced state of subsystem $B$ unchanged. Although measurement breaks the entanglement between two systems—by always mapping $\rho_{AB}$ to a product state—distant parties will observe no correlation between their respective measurement outcomes. In such correlation-free variant of quantum theory, Bell’s inequalities cannot be violated, and entanglement is not available as a resource for distributed tasks.
5.3 Non-collapsing measurements: passive quantum theory
We will now consider measurements that, instead of updating the quantum state, leave it unchanged. [^10] The possibility of non-collapsing measurements was listed by von Neumann [vonneumann_mathematische_1932] as early as 1932 as one of three possible reactions of a physical system to a measurement. Outcomes still occur probabilistically as dictated by the Born rule, yet, in analogy with classical theory, these measurements do not disturb the pre-measurement state. This situation corresponds to a “passive” update rule $\omega^{\mathsf{P}}_{A}$ for non-composite systems,
$\omega^{\mathsf{P}}_{A}\left(P_{x}^{A},\rho_{A}\right)=\text{Tr}\left(P_{x}^{A}\rho_{A}\right)\rho_{A}\,.$ (21)
Non-collapsing measurements, originally introduced in [aaronson_space_2016], cannot be implemented in quantum theory [busch_is_1997]. Indeed, $\omega^{\mathsf{P}}_{A}$ is not convex-linear over $\bar{\mathcal{S}}(\mathcal{H}_{A})$ which implies that, after a passive measurement, one may be able to distinguish between different preparations of the same mixed state. For example, let $\mathcal{H}_{A}=\mathbb{C}^{2}$ and consider the Gemenge
$\mathcal{G}=\{(1/2,|0\rangle\langle 0|),(1/2,|1\rangle\langle 1|)\}\quad\mbox{and}\quad\mathcal{G}^{\prime}=\{(1/2,|+\rangle\langle+|),(1/2,|-\rangle\langle-|)\}\,,$ (22)
where |±$|\pm\rangle=(|0\rangle\pm|1\rangle)/\sqrt{2}$. Both preparations correspond to the maximally mixed state, $\rho_{\mathcal{G}}=\rho_{\mathcal{G}^{\prime}}=\mathbb{I}_{2}/2$. According to the Gemenge-update rule (14) in Sec. 4.3, a passive measurement with outcome $P_{x}^{A}=|0\rangle\langle 0|$ will produce two different sub-normalised density operators,
$\rho_{\mathcal{G}_{x}}=\frac{1}{2}|0\rangle\langle 0|\quad\neq\quad\rho_{\mathcal{G}^{\prime}_{x}}=\frac{\mathbb{I}_{2}}{4}\,.$ (23)
Therefore, the non-collapsing update rule $\omega^{\mathsf{P}}_{A}$ does not satisfy preparation indistinguishability, indicating that density operators do not provide complete descriptions of individual systems.
While complete positivity can be defined for both linear and nonlinear transformations [ando_non-linear_1986], it has been described as “physically unfitting” for nonlinear dynamics [czachor_complete_1998]. In fact, although $\omega^{\mathsf{P}}_{A}$ (or, more precisely, extensions of $\omega^{\mathsf{P}}_{A}$ to $\mathcal{L}(\mathcal{H}_{A})$) is not completely positive, the update rule can still be consistently extended to composite systems. Invoking compositional compatibility (Def. 3), we set
$\omega^{\mathsf{P}}_{AB}\left(P_{x}^{A},\rho_{AB}\right)=\omega^{\mathsf{P}}_{AB}\left(P_{x}^{A}\otimes\mathbb{I}_{B},\rho_{AB}\right)=\text{Tr}\left(P_{x}^{A}\otimes\mathbb{I}_{B}\rho_{AB}\right)\rho_{AB}\,.$ (24)
The map $\omega^{\mathsf{P}}$ satisfies all the conditions in Def. 1 and defines a GUT in which measurements do not modify the state of either the measured system or of the larger composite system to which it belongs. We called this theory passive quantum theory (pQT). An overview of pQT and its key properties has been provided in [fiorentino_quantum_2023].
The absence of any measurement-induced state update allows, in principle, for tomographic reconstruction of the state of an individual system. Thus, the cloning of arbitrary states becomes possible and makes computation in pQT more efficient (in some sense) than in standard quantum theory [busch_is_1997, kent_nonlinearity_2005]. The passive update rule of pQT also satisfies ideality but is not locally tomographic and does not allow for violations of Bell’s inequalities (cf. Table 1).
5.4 Depolarising measurements
Next, consider a theory in which measurements disturb the system to the point that the post-measurement state contains no information about the original state. Repeating measurements on the same system, therefore, extracts no additional information about it.
Updating the pre-measurement state to the maximally mixed state, regardless of the observed outcome, provides one possible realisation of such a theory,
$\omega^{\mathsf{dep}}_{A}\left(P_{x}^{A},\rho_{A}\right)=\text{Tr}\left(P_{x}^{A}\rho_{A}\right)\frac{\mathbb{I}_{A}}{d_{A}}\,,$ (25)
where $d_{A}=\text{dim}(\mathcal{H}_{A})$. The map in Eq. (25) is both convex-linear over $\bar{\mathcal{S}}(\mathcal{H}_{A})$ and completely positive. The set $\{\omega^{\mathsf{dep}}_{A}(P_{x}^{A})\}_{x}$ represents, in fact, a quantum instrument [^11] Specifically, it belongs to the class of trivial instruments. compatible with the observable $\{P_{x}^{A}\}_{x}$.
Following Eq. (18), by defining $\omega^{\mathsf{dep}}_{A,x}(\cdot)\equiv\omega^{\mathsf{dep}}_{A}(P^{A}_{x},\cdot)$, we can extend $\omega^{\mathsf{dep}}_{A}$ to composite systems via
$\omega^{\mathsf{dep}}_{AB}\left(P_{x}^{A},\rho_{AB}\right)=\left(\omega^{\mathsf{dep}}_{A,x}\otimes\mathcal{I}_{B}\right)\left(\rho_{AB}\right)\,.$ (26)
In contrast to the Lüders rule, the depolarising rule $\omega^{\mathsf{dep}}$ does not satisfy composition compatibility. To see this, consider $\rho_{AB}=|00\rangle\langle 00|\in\mathcal{S}(\mathcal{H}_{2}\otimes\mathcal{H}_{2})$ and $P_{x}^{A}=|0\rangle\langle 0|$. We find that
$\omega^{\mathsf{dep}}_{AB}\left(|0\rangle\langle 0|,|00\rangle\langle 00|\right)=\frac{\mathbb{I}_{2}}{2}\otimes|0\rangle\langle 0|\quad\neq\quad\frac{\mathbb{I}_{2}}{2}\otimes\frac{\mathbb{I}_{2}}{2}=\omega^{\mathsf{dep}}_{AB}\left(|0\rangle\langle 0|\otimes\mathbb{I}_{B},|00\rangle\langle 00|\right)\,.$ (27)
This example also demonstrates how composition compatibility could not have been used to extend $\omega^{\mathsf{dep}}_{A}$ to composite systems. The right-hand side of Eq. (27) cannot represent the post-measurement state of the composite system following a local measurement on a single qubit, as it would violate the no-signalling condition A5. See Table 1 for further properties of depolarising measurements.
5.5 Probability-amplifying projective measurements
Suppose an observable is measured twice on the same system in quick succession. Being deterministically repeatable, quantum theory predicts that we will observe the same outcome for the second measurement with certainty. We can construct update rules that satisfy a weaker notion of repeatability, called weak repeatability: the probability of observing an outcome for the second time is greater than the probability of observing it for the first time. Consider the following one-parameter family of update rules for non-composite systems:
$\omega^{\lambda}_{A}\left(P_{x}^{A},\rho_{A}\right)=\text{Tr}{\left(P_{x}^{A}\rho_{A}\right)}\frac{G_{\lambda}^{1/2}\left(P_{x}^{A}\right)\,\rho_{A}\,G_{\lambda}^{1/2}\left(P_{x}^{A}\right)}{\text{Tr}{\left(G_{\lambda}\left(P_{x}^{A}\right)\,\rho_{A}\right)}}\,,$ (28)
where
$G_{\lambda}\left(P_{x}^{A}\right)=\left(1-\lambda\right)P_{x}^{A}+\lambda\mathbb{I}_{A}\,,\quad\lambda\in[0,1]\,.$ (29)
If a measurement outcome $P_{X}^{A}=\sum_{i\in X}|i\rangle\langle i|$ for $X\subseteq\{0,...,d_{A}-1\}$ is obtained given the pure state $|\psi\rangle=\sum_{i=0}^{d_{A}-1}c_{i}|i\rangle\in\mathcal{H}_{A}$, the resulting post-measurement state reads
$\omega^{\lambda}_{A}\left(P_{X}^{A},|\psi\rangle\langle\psi|\right)=\text{Tr}{\left(P_{x}^{A}\rho_{A}\right)}|\psi_{X}\rangle\langle\psi_{X}|\,,$ (30)
where
$|\psi_{X}\rangle=\frac{1}{N}\left(\sum_{j\in X}c_{j}|j\rangle+\sqrt{\lambda}\sum_{i\notin X}c_{i}|i\rangle\right)\,,$ (31)
with $N$ being the normalisation factor. Therefore, the update rule $\omega^{\lambda}_{A}$ projects $|\psi\rangle$ to the pure state $|\psi_{X}\rangle$ and the parameter $\lambda$ determines how likely a second measurement will yield the same outcome as the first, with a lower value of $\lambda$ corresponding to a higher probability of repeating the first outcome. The Lüders rule $\omega^{\mathsf{L}}_{A}$ and the passive rule $\omega^{\mathsf{P}}_{A}$ for non-composite systems are recovered by setting $\lambda=0$ and $\lambda=1$, respectively. Thus, for non-composite systems, the update rule $\omega^{\lambda}_{A}$ indeed smoothly interpolates between quantum theory and pQT. Weak repeatability occurs for $\lambda\in[0,1)$. The map in Eq. (28) is not convex-linear over $\bar{\mathcal{S}}(\mathcal{H}_{A})$ for $\lambda\neq 0$, hence the corresponding theories will not satisfy preparation indistinguishability.
Following Eq. (19), we find that a valid extension of Eq. (28) to composite systems is provided by
$\omega_{AB}^{\lambda}\left(P_{x}^{A},\rho_{AB}\right)=\frac{\omega_{A}^{\lambda}\left(P_{x}^{A},\text{Tr}_{B}\left(\rho_{AB}\right)\right)\otimes\text{Tr}_{A}\left(\rho_{AB}\right)}{\text{Tr}\left(\rho_{AB}\right)}\,.$ (32)
It can be shown that Eq. (32) satisfies all conditions outlined in Def. 1. For each $\lambda\in(0,1)$, the probability-amplifying projective rule $\omega^{\lambda}$ defines a weakly repeatable GUT that is manifestly different from quantum theory. [^12] Setting $\lambda=0$ recovers the locally-Lüders rule $\omega_{AB}^{\mathsf{locL}}$ of Eq. (19). While alternative extensions of $\omega^{\lambda}_{A}$ may in principle be constructed, none would satisfy composition compatibility (Def. 3). Similarly to the depolarising rule $\omega^{\mathsf{dep}}$, it is easy to show that the extension provided by Eq. (16) violates no-signalling A5, as well as local commutativity A6.
5.6 Invalid state-update rules
It is instructive to consider examples of state-update rules that violate at least one of the basic requirements A1-A6 of Def. 1 in Sec. 4.2. We have, in fact, already acknowledged two of them when presenting depolarising measurements (Sec. 5.4) and probability-amplifying projective measurements (Sec. 5.5). Extending the non-composite rules $\omega^{\mathsf{dep}}_{A}$ (25) and $\omega^{\lambda}_{A}$ (28) via compositional compatibility (Def. 3) leads to maps assigning post-measurement states to composite systems in violation of the (quantum) no-signalling principle A5.
Table 2 summarizes the properties of the update rules we will consider in this section.
$\tilde{\omega}^{\mathsf{dep}}$ $\tilde{\omega}^{\lambda}$ $\omega^{\mu}$ $\omega^{\mathsf{vN}}$ $\omega^{\mathsf{U}}$
A1: completeness
A2: context-independence $\times$
A3: local covariance $\times$
A4: self-consistency
A5: no-signalling $\times$ $\times$
A6: local commutativity $\times$ $\times$
Table 2: Summary of the state-updates described in Sec. 5.6 that do not define an update rule as per Def. 1. For each rule, we record satisfied/violated assumptions using ticks and crosses, respectively. The first two, $\tilde{\omega}^{\mathsf{dep}}$ and $\tilde{\omega}^{\lambda}$, denote the extensions to composite systems of $\omega^{\mathsf{dep}}_{A}$ (25) and $\omega^{\lambda}_{A}$ (28), respectively, obtained via composition compatibility (Def. 3).
Consider now a convex mixture of the Lüders $\omega^{\mathsf{L}}$ and the passive $\omega^{\mathsf{P}}$ update rules for composite systems,
$\omega^{\mu}_{AB}\left(P_{x}^{A},\rho_{AB}\right)=\left(1-\mu\right)\omega^{\mathsf{L}}_{AB}\left(P_{x}^{A},\rho_{AB}\right)+\mu\,\omega^{\mathsf{P}}_{AB}\left(P_{x}^{A},\rho_{AB}\right),\quad\quad\mu\in(0,1)\,,$ (33)
where the right-hand side describes a proper mixture. Eq. (33) describes a local measurement that, with some a priori probability $(1-\mu)$, collapses the state of the composite system as prescribed by quantum theory, or, with probability $\mu$, leaves the state unchanged. For $\mu\in(0,1)$, the probability of observing outcome $P^{A}_{x}$ in a possible second measurement is, on average, higher than in the first measurement. In other words, (33) provides an alternative way to satisfy weak repeatability, different from the probability-amplifying projections $\omega^{\lambda}$ introduced in Sec. 5.5.
The “real” post-measurement state of the system is either $\omega^{\mathsf{L}}_{AB}\left(P_{x}^{A},\rho_{AB}\right)$ or $\omega^{\mathsf{P}}_{AB}\left(P_{x}^{A},\rho_{AB}\right)$; one can imagine that this is determined by the value of some hidden variable $\lambda\in\{0,1\}$. The probability distribution of this hidden variable is the same across all experimental runs—i.e. $\text{prob}(\lambda=0)=1-\mu$, and $\text{prob}(\lambda=1)=\mu$—ensuring that it is independent of the context of a particular experiment. This guarantees that $\omega^{\mu}$ satisfies context-independence A2. Moreover, (33) can be shown to satisfy completeness A1, local covariance A3, self-consistency A4 and no-signalling A5, since $\omega^{\mu}$ is a convex mixture of mappings that individually satisfy these conditions. Surprisingly, however, we find that (33) violates local commutativity A6, despite both $\omega^{\mathsf{L}}$ and $\omega^{\mathsf{P}}$ satisfying it.
The violation can be seen in a simple example. Let $\rho_{AB}=|\Phi^{+}\rangle\langle\Phi^{+}|\in\mathcal{S}(\mathcal{H}_{2}\otimes\mathcal{H}_{2})$, where $|\Phi^{+}\rangle=(|00\rangle+|11\rangle)/\sqrt{2}$ is a maximally entangled state, and consider $P_{x}^{A}=|0\rangle\langle 0|$, and $P_{y}^{B}=|1\rangle\langle 1|$. Recall that $\omega^{\mu}$ maps pure states to proper mixtures; hence, when representing sequential measurements, the second map must be applied to each pure state in the corresponding Gemenge, in line with the discussion in Sec. 4.3. If $P^{A}_{x}$ is obtained in a local measurement on $\mathcal{H}_{A}$ before the local measurement of $\mathcal{H}_{B}$ yields $P^{B}_{y}$, the state of the composite system updates as
$|\Phi^{+}\rangle\langle\Phi^{+}|\,\stackrel{{\scriptstyle(x\prec y)}}{{\mapstochar\relbar\joinrel\relbar\joinrel\relbar\joinrel\rightarrow}}\,\frac{\mu}{4}\left(\left(1-\mu\right)|11\rangle\langle 11|+\mu|\Phi^{+}\rangle\langle\Phi^{+}|\right)\,.$ (34)
If instead the measurement yielding $P^{B}_{y}$ precedes the one yielding $P^{A}_{x}$, a different update occurs,
$|\Phi^{+}\rangle\langle\Phi^{+}|\,\stackrel{{\scriptstyle(y\prec x)}}{{\mapstochar\relbar\joinrel\relbar\joinrel\relbar\joinrel\rightarrow}}\,\frac{\mu}{4}\left(\left(1-\mu\right)|00\rangle\langle 00|+\mu|\Phi^{+}\rangle\langle\Phi^{+}|\right)\,.$ (35)
Since the final state depends on the order in which the local measurements are performed, Eq. (33) violates local commutativity A6. As a result, correlations between measurements on different subsystems cannot be unambiguously determined.
The projection postulate introduced by von Neumann [vonneumann_mathematische_1932] is a notable example of state-update that does not define a valid update rule. For non-degenerate measurements, von Neumann’s and Lüders’ postulates coincide. However, according to von Neumann, outcome degeneracy of quantum measurements—represented by projectors of rank greater than one—arises from classical post-processing of non-degenerate measurements. That is, one does not directly implement a measurement of a degenerate observable $M$; instead, a measurement of a refinement of $M$, i.e. a non-degenerate observable $M^{\prime}$ commuting with $M$, is carried out, and the outcomes are coarse-grained. Given a degenerate observable, there exist, however, infinitely many refinements. Consequently, the post-measurement state will not only depend solely on the pre-measurement state $\rho_{A}$ (which we can assume to be pure, without loss of generality) and the observed degenerate outcome $P^{A}_{x}$, but also on the specific refinement chosen. More precisely, letting $\delta=\text{Tr}{(P^{A}_{x})}$ denote the “cardinality” of the degeneracy of outcome $P^{A}_{x}$, the updated state depends on the subset $\mathcal{B}_{x}=\{P^{A}_{x_{i}}\}_{i=1}^{\delta}\subseteq M^{\prime}$ spanning the degenerate subspace $\mathcal{H}_{x}=P^{A}_{x}\mathcal{H}$. The state-update for non-composite systems thus obeys
$\omega^{\mathsf{vN}}_{A}\left(P^{A}_{x},\rho_{A},\mathcal{B}_{x}\right)=\sum_{i=1}^{\delta}\omega^{\mathsf{L}}_{A}\left(P^{A}_{x_{i}},\rho_{A}\right)\,,$ (36)
where the right-hand side describes a proper mixture. As a result, von Neumann’s postulate—despite being, like Lüders’ original postulate ($\mathsf{L}$), concerned only with the state-update of the measured system—violates context-independence A2 and, therefore, does not give rise to a valid update rule. [^13] One could, however, modify (36) into a valid update rule by, for example, fixing a preferred set $\mathcal{B}_{x}$ for each experiment yielding the outcome $P^{A}_{x}$, and subsequently extend the map to composite systems via complete positivity.
Finally, we present a simple example of an update rule that violates local covariance A4. Assume that any measurement on a system causes the pre-measurement state to undergo a fixed, observer-independent and non-trivial [^14] If $U=\mathbb{I}$, the passive measurements of Sec. 5.3 result consistent with A4. unitary transformation $U\in\mathcal{U}(\mathcal{H})$,
$\omega^{\mathsf{U}}_{AB}\left(P^{A}_{x},\,\rho_{AB}\right)=\text{Tr}\left(P_{x}^{A}\text{Tr}_{B}\left(\rho_{AB}\right)\right)\,\left(U\otimes\mathbb{I}_{B}\right)\,\rho_{AB}\,\left(U^{\dagger}\otimes\mathbb{I}_{B}\right)\,.$ (37)
The rule $\omega^{\mathsf{U}}_{AB}$ is not locally covariant when using local unitaries $U_{A}$ that do not commute with $U$: different observers would provide inconsistent descriptions of the same measurement.
6 How to recover the projection postulate
We have set up a framework of hypothetical update rules that do give rise to consistent foils of quantum theory, and we explored their properties. It is natural to reverse our approach by turning to the question of singling out the Lüders rule of quantum theory among the set of valid updates. The correlation-free version of quantum theory (Sec. 5.2)—as well as any suitable modification of von Neumann’s projection (36)—shows that the Lüders rule cannot be recovered uniquely from assuming deterministic repeatability alone. Similarly, passive quantum theory (Sec. 5.3) provides another example of an update rule satisfying ideality.
Lüders presents two main objections to von Neumann’s projection postulate [luders_uber_1950]. Firstly, he argues that the post-measurement state should depend only on the outcome and the initial state. This requirement is equivalent to context-independence A2. Secondly, he suggests that
“the measurement of a highly degenerate quantity permits only relatively weak assertions regarding the considered ensemble. For that reason, the resulting change in state should likewise be small” ([luders_concerning_2006], p. 665).
In the limit of the trivial observable represented by the identity operator $\mathbb{I}$, a measurement reveals no information about the original state and it should have no effect on it. The quoted property describes a form of trade-off between the applied disturbance caused by a measurement and the information acquired from the outcome. The following operational property for non-composite systems, which we term coherence, implies the desired behaviour for measurements of the trivial observable.
Definition 4 (Coherence).
An update rule for non-composite systems $\omega_{A}$ satisfies coherence if the outcome probability distribution of any observable $M$ is not affected by a prior measurement of a coarse-graining of $M$ (i.e. an observable for which $M$ is a refinement). Mathematically, for every $P^{A}_{x},P^{A}_{y}\in\mathcal{P}(\mathcal{H}_{A})$ such that $P^{A}_{x}\,P^{A}_{y}=P^{A}_{y}\,P^{A}_{x}=P^{A}_{x}$, and every state $\rho_{A}\in\bar{\mathcal{S}}(\mathcal{H}_{A})$, we must have
$\text{Tr}{\left[P^{A}_{x}\,\omega_{A}\left(P^{A}_{y},\rho_{A}\right)\right]}=\text{Tr}{\left(P^{A}_{x}\rho_{A}\right)}\,.$ (38)
Consider an infinite ensemble prepared in an arbitrary quantum state, and examine the following two scenarios: $(i)$ a measurement of $M$ is performed on each element of the ensemble; $(ii)$ a measurement of a coarse-graining of $M$ followed by a measurement of $M$ are performed on each element of the ensemble. The property of coherence stipulates that the outcome probabilities observed when measuring $M$ in scenarios $(i)$ and $(ii)$ must be identical. In fact, Eq. (38) ensures that, for all quantum states and outcomes $P^{A}_{x}$ and $P^{A}_{y}$ such that $P^{A}_{x}\,P^{A}_{y}=P^{A}_{y}\,P^{A}_{x}=P^{A}_{x}$, the relation
$\text{prob}(o_{1}=P^{A}_{y})\,\text{prob}(o_{2}=P^{A}_{x}|o_{1}=P^{A}_{y})=\text{prob}(o_{1}=P^{A}_{x})$ (39)
is satisfied, where $o_{1}$ and $o_{2}$ denote the first and second outcomes of two consecutive measurements, respectively. Here, $P^{A}_{x}$ denotes the outcome of the $M$ measurement, while $P^{A}_{y}$ denotes the outcome of the coarse-graining. If $P^{A}_{x}\neq P^{A}_{y}$, then $\mathcal{H}_{x}\subset\mathcal{H}_{y}$ (where $\mathcal{H}_{i}=P_{i}^{A}\mathcal{H}$), meaning that $P^{A}_{y}$ provides less information about the original state than $P^{A}_{x}$. As a result, the effect on the state cannot be arbitrarily strong but must “preserve coherence” in the subspace $\mathcal{H}_{y}$, in the sense that $\text{prob}(o_{2}=P^{A}_{x}|o_{1}=P^{A}_{y})=\text{const.}\times\text{prob}(P^{A}_{x})$ for all $P^{A}_{x}\in\mathcal{P}(\mathcal{H}_{y})$. In the limiting case of $P^{A}_{y}=\mathbb{I}_{A}$, one obtains $\omega_{A}\left(\mathbb{I}_{A},\rho_{A}\right)\propto\rho_{A}$, in agreement with Lüders’ remark.
In some form or another, the assumption of coherence has appeared already in the literature, both in the context of quantum theory [bell_moral_1966, herbut_compatibility_2004, khrennikov_neumann_2009] and of generalised probabilistic theories [kleinmann_sequences_2014]. For measurements on non-composite quantum systems, coherence entails both deterministic repeatability [kleinmann_sequences_2014] and ideality.
Lemma 1.
A coherent update rule for non-composite systems $\omega_{A}$ satisfies both deterministic repeatability and ideality.
Proof.
Consider an arbitrary projector $P_{x}^{A}$ and a system residing in some state $\rho_{A}\in\mathcal{S}(\mathcal{H}_{A})$ such that $\text{Tr}(P_{x}^{A}\rho)\neq 0$. Letting $P_{y}^{A}=P_{x}^{A}$ in Eq. (38) leads to $\text{Tr}\left(P_{x}^{A}\,\omega_{A}\left(P_{x}^{A},\rho_{A}\right)\right)=\text{Tr}\left(P_{x}^{A}\rho_{A}\right)$, which implies that, if repeated, the measurement will yield the same outcome with probability $1$. In other words, $\omega_{A}\left(P_{x}^{A},\rho_{A}\right)$ has support in the subspace $\mathcal{H}_{x}=P_{x}^{A}\mathcal{H}_{A}$ only, i.e. $\omega_{A}\left(P_{x}^{A},\rho_{A}\right)\in\bar{\mathcal{S}}(\mathcal{H}_{x})$. Therefore, $\omega_{A}$ satisfies deterministic repeatability.
Following the same argument, we know that $\omega_{A}\left(P_{y}^{A},\rho_{A}\right)$ has support in $\mathcal{H}_{y}$ only. Let the initial state $\rho_{A}$ also have support in $\mathcal{H}_{y}$ only, i.e. $\text{Tr}(P_{y}^{A}\rho_{A})=1$, and consider an informationally complete set of observables on $\mathcal{H}_{y}$. If $\omega_{A}\left(P_{y}^{A},\rho_{A}\right)\neq\rho_{A}$, then for at least one of the outcomes of an observable in the set, say $P_{x}^{A}\in\mathcal{P}(\mathcal{H}_{y})$ (which must satisfy $P^{A}_{x}\,P^{A}_{y}=P^{A}_{x}$), we have $\text{Tr}{\left(P_{x}^{A}\omega_{A}\left(P_{y}^{A},\rho_{A}\right)\right)}\neq\text{Tr}{\left(P_{x}^{A}\rho_{A}\right)}$, in contradiction with (38). Therefore, the update rule does not disturb the states of measured systems when the outcome is certain, i.e. $\omega_{A}$ satisfies ideality. ∎
Indeed, coherence alone is sufficient to derive the single-system Lüders rule $\omega^{\mathsf{L}}_{A}$, i.e. the original projection postulate ($\mathsf{L}$). This argument was first presented in 1966 [bell_moral_1966] and similar versions of it have since been rediscovered several times; see Appendix A for a summary. Below, we present a simple proof of this result.
Theorem 1.
For non-composite systems, the Lüders rule $\omega^{\mathsf{L}}_{A}$ is the only coherent update rule.
Proof.
Substituting $\omega^{\mathsf{L}}_{A}$ in (38) leads to
$\text{Tr}{\left(P^{A}_{x}P^{A}_{y}\rho_{A}\,P^{A}_{y}\right)}=\text{Tr}{\left(P^{A}_{x}\,\rho_{A}\right)}\,,$ (40)
which holds for all $\rho_{A}$ and $P^{A}_{x}\,P^{A}_{y}=P^{A}_{y}\,P^{A}_{x}=P^{A}_{x}$. Hence the Lüders projection for single systems is coherent.
To show the converse, suppose $\omega_{A}$ satisfies coherence but $\omega_{A}\neq\omega^{\mathsf{L}}_{A}$. Then, there must exist some outcome $P^{A}_{y}\in\mathcal{P}(\mathcal{H}_{A})$ and state $\rho_{A}\in\mathcal{S}(\mathcal{H}_{A})$ such that $\omega_{A}\left(P^{A}_{y},\rho_{A}\right)\neq\omega^{\mathsf{L}}_{A}\left(P^{A}_{y},\rho_{A}\right)$. Since both rules are coherent, hence deterministically repeatable by Lemma 1, both updated states must lie within the same subspace defined by $P^{A}_{y}$, i.e. $\omega_{A}\left(P^{A}_{y},\rho_{A}\right)$, $\omega^{\mathsf{L}}_{A}\left(P^{A}_{y},\rho_{A}\right)\in\bar{\mathcal{S}}(\mathcal{H}_{y})$. Consider a set of observables of $\mathcal{H}_{y}$ which allow one to reconstruct any quantum state in $\bar{\mathcal{S}}(\mathcal{H}_{y})$ tomographically. The states $\omega_{A}\left(P^{A}_{y},\rho_{A}\right)$ and $\omega^{\mathsf{L}}_{A}\left(P^{A}_{y},\rho_{A}\right)$ must yield different probabilities for at least one outcome $P^{A}_{x}\in\mathcal{P}(\mathcal{H}_{y})$ in the set,
$\text{Tr}{\left[P^{A}_{x}\,\omega_{A}\left(P^{A}_{y},\rho_{A}\right)\right]}\neq\text{Tr}{\left[P^{A}_{x}\,\omega^{\mathsf{L}}_{A}\left(P^{A}_{y},\rho_{A}\right)\right]}\,.$ (41)
Substituting $\omega^{\mathsf{L}}_{A}\left(P^{A}_{y},\rho_{A}\right)=P^{A}_{y}\,\rho_{A}\,P^{A}_{y}$ and using the fact that $P^{A}_{x}P^{A}_{y}=P^{A}_{y}\,P^{A}_{x}=P^{A}_{x}$, we obtain
$\text{Tr}{\left[P^{A}_{x}\,\omega_{A}\left(P^{A}_{y},\rho_{A}\right)\right]}\neq\text{Tr}{\left(P^{A}_{x}\rho_{A}\right)}\,,$ (42)
which contradicts the assumption of coherence. We thus conclude that $\omega_{A}=\omega^{\mathsf{L}}_{A}$. ∎
The information-disturbance trade-off principle, as encapsulated by the property of coherence, completely characterises the update rule for quantum measurements on non-composite systems. However, there do exist GUTs with different update rules for multi-partite systems that reduce to the single-system Lüders rule. The correlation-free theory of Sec. 5.2 defined by the update rule $\omega^{\mathsf{locL}}$ is an explicit example. Straightforward generalisations of coherence to composite systems do not seem to be sufficiently strong to generalise the proof of Theorem 1. For example, the assumption that a sequence of measurements of $P_{y}^{A}$ and $P_{x}^{A}$ (see Def. 4) on subsystem $\mathcal{H}_{A}$ leads to the same post-measurement state for $\mathcal{H}_{AB}$ as a single measurement of $P_{x}^{A}$ fails to rule out correlation-free quantum theory.
Additional assumptions are necessary to fully recover the Lüders rule $\omega^{\mathsf{L}}_{AB}$ for composite systems. One such assumption is composition compatibility (Def. 3)—see Theorem 2 below. This final step in deriving the projection postulate is often overlooked [bell_moral_1966, herbut_compatibility_2004, martinez_search_1990], as the extension of $\omega^{\mathsf{L}}_{A}$ to local measurements on composite systems is typically assumed implicitly, despite not being stated in or implied by the standard postulates.
Theorem 2.
The Lüders rule $\omega^{\mathsf{L}}$ is the unique generalised state-update rule that satisfies both coherence and composition compatibility.
Proof.
Theorem 1 derives the Lüders projection $\omega^{\mathsf{L}}_{A}$ for non-composite systems from coherence alone. Then, the quantum-mechanical extension to composite systems $\omega^{\mathsf{L}}_{AB}$ is fixed by Eq. (16) defining composition compatibility, as already discussed in Sec. 5.1. ∎
This derivation of the Lüders rule does not rely on assuming linearity of the update rule. There may, of course, be other physically appealing operational principles ruling out all update rules but the Lüders rule. For instance, it seems promising to explore the implications of nonlocality. Could quantum theory be singled out by imposing specific constraints on the degree of nonlocality exhibited by generalised state-update theories?
7 Conclusions
7.1 Summary
We have introduced a framework that allows us to systematically investigate modifications of the projection postulate. Theories with generalised state-update rules (GURs) differ from quantum theory only in assigning other post-measurement states.
Upon formalising the concept of state-update rules, we notice that the quantum mechanical Lüders rule is typically presented in an operationally incomplete way. Spelling out the projection postulate (1) for single systems does not entail a unique update rule when local measurements are performed on composite systems. In standard quantum theory, the necessary extension from single to composite systems is normally assumed implicitly to have a specific form. This observation is a simple example of the subtleties in the definition of update rules. The specific update rule used in composite quantum systems is, of course, well motivated by experimentally verified correlations between measurement outcomes obtained for entangled states.
Any valid state-update rule must satisfy six core requirements which include the uniqueness of post-measurement states, no-signalling, and consistency for measurements carried out by observers with access to subsystems only. We have shown that quantum theory and passive quantum theory [fiorentino_quantum_2023] satisfy the requirements. Three new examples of consistent generalised state-update rules have been defined, each one leading to a toy theory structurally different from quantum theory.
Correlation-free quantum theory turns out to be locally indistinguishable from standard quantum theory since the update rule for single systems coincides with the projection postulate. However, the proposed update rule does not extend in the standard way to composite systems. The update rule is “locally Lüders” but local measurements performed on entangled states do not create the correlations known from quantum theory.
Depolarising measurements are described by an update rule that sends any pre-measurement state to the maximally mixed state, irrespective of the measurement outcome. The update effectively acts as a completely depolarising channel, flattening the probability distribution for subsequent outcomes of measurements. No information about the pre-measurement state is retained.
Probability-amplifying projective rules lead to post-measurement states that assign a larger probability to the measurement outcome than the pre-measurement state did. The standard projective update rule is the extreme case of this rule: the outcome of the second measurement is necessarily the same as that of the first one, implying deterministic repeatability. The Lüders rule for single systems can be approximated arbitrarily well by probability-amplifying projective state updates.
To avoid signalling in theories with depolarising or probability-amplifying update rules, careful adjustments are needed when extending them to composite systems. The ways in which state-update rules may violate core requirements have been illustrated by a number of (ultimately invalid) GURs.
Having established a variety of acceptable update rules, it is natural to look for properties that single out the Lüders rule. We have shown that the property of coherence uniquely implies the Lüders rule for non-composite systems. This result agrees with earlier ones based on similar concepts describing a form of information-disturbance trade-off. However, our systematic approach to GURs highlights the fact that known derivations of the Lüders rule are incomplete since they typically do not take composite systems into account. We find that coherence on its own is insufficient to single out the projection postulate when considering composite systems; equally, composition compatibility alone is not strong enough to do so. However, combining it with coherence implies the projective Lüders rule for composite systems, too (Theorem 2).
7.2 Discussion and outlook
The framework of generalised state-update rules we present affords us with a bird’s-eye view of the projection postulate and physically reasonable modifications thereof, their basic features and their implications. It is useful to think of update rules as generalisations of quantum instruments that are neither linear nor completely positive, all the while still providing consistent descriptions of measurements on (hypothetical) quantum systems.
Non-linear update rules may enable observers to distinguish between proper and improper mixtures. In the corresponding foil theories, density operators will not provide a complete description of the hypothetical physical systems. Importantly, complete positivity is not strong enough to ensure that such measurement-induced transformations assign post-measurement states to composite systems consistently. In the words of [czachor_complete_1998], complete positivity is “physically unfitting” for this purpose.
Reconstructions of quantum theory that engage—either directly or indirectly—with post-measurement states usually assume measurement-induced state transformations to be linear in the pre-measurement state [barrett_information_2007, kleinmann_sequences_2014, masanes_measurement_2019]. This assumption, often left implicit [stacey_masanes-galley-muller_2022, kent_measurements_2023], excludes a priori certain types of measurement behaviours—such as the trivial non-collapsing rule of “passive quantum theory”—that are, in fact, operationally well-defined and consistent with fundamental principles such as no-signalling (see Sec. 4.2). Preliminary results based on the framework of GURs (as well as earlier works, e.g. [flatt_gleason-busch_2017]) suggest that the linearity of update rules could be derived from other, more natural operational assumptions.
It seems promising to investigate the role of nonlocality from the perspective of theories with generalised update rules. Do other update rules exist that are capable of replicating the correlations found in quantum theory? What is more, can they lead to super-quantum correlations?
Acknowledgements
The authors are grateful for support through the Leverhulme grant RPG-2024-201.
Appendix A Summary of derivations of the Lüders rule
Derivations of the Lüders rule have a long history, with similar arguments being made over the years. We will briefly review derivations known to us, highlighting the assumptions made. A key observation is that most of the justifications focus on non-composite systems, either ignoring measurements on composite systems or making implicit assumptions about them.
In 1966, Bell and Nauenberg showed that the quantum mechanical state-update can be obtained by assuming that the outcome probability distribution of a common refinement for commuting observables $M_{1}$ and $M_{2}$ equals the joint probability distribution obtained by measuring $M_{1}$ and $M_{2}$ sequentially, in any order [bell_moral_1966]. This assumption is mathematically equivalent to that of coherence (Def. 4) employed in the proof of Theorem 1 in Sec. 6. To the best of our knowledge, the only reference to this early operational derivation of the Lüders rule in non-composite systems is by Herbut [herbut_compatibility_2004]. Over the last 40 years, the argument by Bell and Nauenberg has been rediscovered at least three times, using the concept of coherence.
In 1983, Cassinelli and Zanghì formulated quantum theory in terms of a generalised probability space and showed how to recover the single-system Lüders rule in this context. To do so, they describe the state-update following measurements in terms of a generalised conditional probability defined on the generalised probability space—i.e. a probability measure on $\mathcal{P}(\mathcal{H})$ compatible with Born’s rule and satisfying “coherence” [cassinelli_conditional_1983].
In 2009, Khrennikov showed that the Lüders rule for single systems follows from von Neumann’s projection postulate for non-degenerate observables if “coherence” is assumed [khrennikov_neumann_2009]. Claiming that “this important observation remained unnoticed for the past 70 years”, the author appears to be unaware of Bell and Nauenberg’s work.
In 2014, Kleinmann developed a formalism for sequential measurements in a broad class of generalised probabilistic theories (GPTs), introducing a generalisation of the Lüders rule based on coherence [kleinmann_sequences_2014]. A simple derivation of the non-composite Lüders rule based on coherence is also given, without reference to earlier work. When applied to quantum theory, i.e. in a Hilbert-space setting for states, Kleinmann’s approach reproduces convex-linear update rules for single systems. This restriction is due to assuming preparation indistinguishability throughout (see Sec. 5.1), thereby excluding operationally valid alternatives such as the passive update rule from the outset.
Herbut actually presents two distinct derivations applying to measurements on non-composite systems in a paper from 1969 [herbut_derivation_1969]. In an information-theoretic approach, he postulates (i) deterministic repeatability (see Sec. 2.1) and that (ii) the post-measurement state minimises its distance from the pre-measurement state. The Lüders rule is then shown to map the pre-measurement state to the least distinguishable state compatible with outcome repeatability. This result remained largely unnoticed, while anticipating later arguments deriving the state-update from minimisation principles, including those by Marchand [marchand_statistical_1977], Dieks [dieks_distance_1983], and Hadjisavvas [hadjisavvas_distance_1981]. More recently, the single-system Lüders rule was obtained from minimising non-symmetric quantum information distances [kostecki_luders_2014, hellmann_quantum_2016], thereby establishing a mathematical connection to the Bayes-Laplace rule. Ozawa [ozawa_quantum_1984, ozawa_quantum_1998] derives the update rule from Bayes’ rule by assuming the standard joint probability formula for local quantum measurements.
Herbut’s second derivation in [herbut_derivation_1969] takes an operational approach. In addition to deterministic repeatability, it is assumed that performing an unconditional measurement of observable $M$ does not affect the expectation values of any observables commuting with $M$. The argument focuses on measurements on single systems. However, Herbut also briefly considers composite systems, discussing in particular how local Lüders measurements modify joint states without enabling signalling. When doing so, he implicitly adopts the standard extension to composite systems (see Sec. 5.1). What is more, Herbut takes convex-linearity over states to be a fundamental requirement for state updates, without justifying this assumption.
Herbut revisits the question in a 1974 paper [herbut_minimal_1974]. Here, the Lüders rule for single systems is obtained by combining the assumption of deterministic repeatability with the “preservation of sharp values”. An observable $M$ has a sharp value $x$ if $\text{Tr}(P_{x}^{A}\rho)=1$ holds for some $P_{x}\in M$. This value is preserved if, given an initial state $\rho$ with a sharp $M$-value $x$, a measurement of any observable commuting with $M$ leaves unchanged the value $x$, regardless of the observed outcome. When proving this result for improper mixtures, Herbut implicitly adopts “composition compatibility” to describe local measurements.
Martinez [martinez_search_1990] offers an early account of known derivations of the Lüders rule including Herbut’s work but not Bell and Nauenberg’s contribution. The paper also demonstrates that the converse of Lüders theorem [luders_uber_1950, busch_quantum_1996] implies the quantum mechanical collapse for non-composite systems: if, for any pair of commuting observables $M_{1}$ and $M_{2}$ and any initial state, we assume that the sequence of measurements $M_{1}\prec M_{2}\prec M_{1}$ always produces identical outcomes for $M_{1}$ regardless of the outcome of the $M_{2}$ measurement, then the post-measurement state must necessarily be given by the Lüders rule. Martinez further examines a quantum-logical derivation employing the “Sasaki hook” presented in [friedman_quantum_1978]. The extension of the update rule to local measurements on subsystems is not considered.
The proof that establishes the Lüders instrument as the unique ideal quantum instrument for sharp discrete measurements dates back to Davies [davies_quantum_1976], as recognised in [busch_quantum_1996, busch_repeatable_1995]. The notion of a quantum instrument incorporates two key assumptions: convex-linearity of the single-system state-update rule, and the standard linear extension to composite systems enabled by the complete positivity of $\omega_{A}$ (see Eq. (18) in Sec. 5.1). This argument for isolating the Lüders collapse for single systems has been rediscovered occasionally (see  [marinkovic_note_1983], for example).
In 2019, Masanes et al. advance the argument that both the Born rule and the post-measurement state rule can be derived from the other postulates of quantum theory when supplemented by suitable operational constraints [masanes_measurement_2019]. In doing so, they consider alternative rules for computing outcome probabilities, explicitly discussing both composite systems and local measurements. Having established the uniqueness of the Born rule within their framework, they assert that quantum instruments constitute the only consistent description of state updates. This conclusion, along with their broader framework, has attracted some criticism [stacey_masanes-galley-muller_2022, galley_reply_2022, kent_measurements_2023, masanes_response_2023]. The argument for quantum instruments depends on an assumption regarding sequential measurements that, as already shown in [flatt_foundations_2019], is mathematically equivalent to requiring convex-linearity of state updates.
Conceptually, Masanes et al.’s approach mirrors Kleinmann’s framework [kleinmann_sequences_2014] in that it is limited a priori to update rules satisfying preparation indistinguishability. This restriction explains why they adopt the completely positive extension (18) for local measurements on composite systems: any other extension would necessarily violate their assumption of preparation indistinguishability.
In retrospect, the Lüders rule is seen to have been derived on the basis of operational, information-theoretic, quantum-logical, and epistemically motivated assumptions. Most of the derivations are limited to single-system update rules, and their extension to local measurements on subsystems is either ignored or tacitly assumed. The extension of state-update rules is, however, not trivial, since the single-system rule alone does not determine it. Additional structural assumptions such as preparation indistinguishability or composition compatibility (cf. Theorem 2) are required to exclude alternative update rules in the multi-partite case. None of these properties are contained in the postulates ($\mathsf{S}$) to ($\mathsf{B}$) of quantum theory.
[…]
[heading=bibintoc]

### v2

@@
-pQT_AMTs.bib
+Beyond the Projection Postulate and Back: Quantum Theories with Generalised State-Update Rules
 Abstract
 Are there consistent and physically reasonable alternatives to the projection postulate?
-Does it have unique properties compared to acceptable alternatives?
+Does it have unique properties compared with acceptable alternatives?
 We answer these questions by systematically investigating hypothetical state-update rules for quantum systems that Nature could have chosen over the Lüders rule.
@@
 Particular attention will be paid to consistently defining post-measurement states when performing local measurements in composite systems.
-Explicit examples of valid unconventional update rules are presented, each of them resulting in a distinct, well-defined foil of quantum theory.
+Explicit examples of valid unconventional update rules are presented, each resulting in a distinct, well-defined foil of quantum theory.
 This framework of state-update rules allows us to identify operational properties that distinguish the projective update rule from all others and to put earlier derivations of the projection postulate into perspective.
 1 Introduction
-The main goal of this contribution is to identify operational properties that distinguish the quantum-mechanical projection postulate from other physically reasonable, but hypothetical, state-update rules.
-Justifications of Lüders rule have, not surprisingly, a long history, and a recent debate revolves around the claim that it can be derived from the other postulates of quantum theory (cf.
+The main goal of our contribution is to identify operational properties that distinguish the quantum-mechanical projection postulate from hypothetical but physically reasonable state-update rules.
+Not surprisingly, attempts to justify Lüders rule (for which both Dirac and von Neumann deserve credit) [^2] In view of the historic developments, it would be appropriate to speak of the “Dirac–von Neumann-Lüders rule” [undefaay] but we will continue using “Lüders rule” for simplicity.
+However, the projection postulate as formulated by Dirac [undefac] and Lüders [undefaaf] differs from von Neumann’s version [undefaaz] in key aspects; see Sec.
+5.6.
+have a long history, and a recent debate revolves around the claim that it can be derived from the other postulates of quantum theory (cf.
 Sec.
@@
 Many of these approaches have in common that their focus is on non-composite systems, either ignoring measurements on composite systems or making implicit assumptions about their consequences.
-We study the nature of the projection postulate systematically by creating foil theories closely resembling quantum theory but equipped with different generalised state-update rules (GURs).
+Our main tool will be the construction of foil theories of quantum theory (cf.
+Sec.
+2.2), i.e.
+toy theories that help “to highlight the distinctive characteristics of quantum theory by contrasting with it” (adapted from [undefu]).
+Thus, we investigate the nature of the projection postulate by comparing quantum theory with similar theories that are equipped with different generalised state-update rules (GURs).
 In other words, we introduce generalised state-update theories (GUTs) for which the post-measurement states associated with a given outcome are not necessarily those prescribed by the Lüders rule.
@@
 Using quantum systems as information carriers has led to applications without classical equivalents.
-Quantum measurements, in particular, provide a resource that contributes, for example, to secure communication protocols [bennett_quantum_2014], quantum teleportation [bennett_teleporting_1993] device-independent random number generators [colbeck_quantum_2009] and to speedy algorithms when using measurement-based quantum computation [briegel_measurement-based_2009].
-The advantages of quantum measurements arise from fundamental differences when compared to measurements in classical theories.
+Quantum measurements, in particular, provide a resource that contributes, for example, to secure communication protocols [undefk], quantum teleportation [undefl] device-independent random number generators [undefv], and to speedy algorithms when using measurement-based quantum computation [undefo].
+The advantages of quantum measurements arise from fundamental differences when compared with measurements in classical theories.
 Classical measurements are deterministic in that they reveal pre-existing values of observables, and the effect on the state of the observed system is negligible, at least in principle.
 In quantum theory, measurement outcomes are probabilistic, and the state of a system after a measurement coincides with the pre-measurement state only in exceptional cases.
-Phenomenologically, the quantum mechanical state update is governed by the projection postulate, which describes a “collapse” of the state to an eigenstate of the operator representing the observed outcome.
-The collapse, also known as state reduction or Lüders rule [luders_uber_1950], ensures deterministic repeatability: two measurements of an observable always yield identical outcomes when implemented in quick succession [compton_directed_1925].
-The standard postulates of quantum theory neither explain whether the state-update is a physical process [bohm_suggested_1952, dewitt_many-worlds_2015], nor do they define what constitutes a measurement [bell_against_1990, bacciagaluppi_heisenberg_2009].
-The foundational questions resulting from this dilemma are often grouped under the term “measurement problem[busch_quantum_1996, norsen_foundations_2017].
+Phenomenologically, the quantum mechanical state update is governed by the projection postulate, which describes a “collapse” of the state to an eigenstate of the operator representing the observed outcome [undefac, undefaaz, undefaaf].
+The collapse, also known as state reduction or Lüders rule [undefaaf], ensures deterministic repeatability: two measurements of an observable always yield identical outcomes when implemented in quick succession [undefw].
+The standard postulates of quantum theory neither explain whether the state-update is a physical process [undefn, undefaa], nor do they define what constitutes a measurement [undefh, undefc].
+The foundational questions resulting from this dilemma are often grouped under the term “measurement problem[undefr, undefaal].
 The present work does not attempt to resolve the interpretational challenges posed by this long-standing problem.
-Justifications of the Lüders rule have been given [bell_moral_1966, herbut_derivation_1969, hellmann_quantum_2016, kostecki_luders_2014, martinez_search_1990] (see Appendix A for a survey), and the difference between von Neumann’s and Lüders’ formulations of the projection postulate has been discussed repeatedly [vonneumann_mathematische_1932, luders_uber_1950, busch_luders_1998, sudbery_whose_2024].
-Some authors have argued that the projection postulate follows from the other standard postulates of quantum theory [ozawa_quantum_1984, masanes_measurement_2019]; however, these derivations—one of which was the subject of a recent debate [stacey_masanes-galley-muller_2022, galley_reply_2022, kent_measurements_2023, masanes_response_2023]—appear to invariably rely on additional assumptions.
-Occasionally, alternative post-measurement rules have been considered, either in the context of standard quantum theory [kent_nonlinearity_2005, kent_quantum_2021, masanes_measurement_2019] or of generalised probabilistic theories [kleinmann_sequences_2014].
+Derivations of the Lüders rule have been given [undefi, undefas, undefaq, undefaad, undefaai] (see Appendix A for a survey), and the difference between von Neumann’s and Lüders’ formulations of the projection postulate has been discussed repeatedly [undefaaf, undefs, undefaay].
+Some authors have argued that the projection postulate follows from the other standard postulates of quantum theory [undefaan, undefaak] and the debate of a recent derivation is ongoing [undefaaw, undefak, undefax, undefaaj, undefaav].
+Occasionally, alternative post-measurement rules have been considered, either in the context of standard quantum theory [undefay, undefaz, undefaak] or of generalised probabilistic theories [undefaab].
 A systematic investigation, as proposed here, has apparently not been carried out before.
 2.2 Foil theories
-Our approach is inspired by the constructive role of foil theories for the understanding of quantum theory.
+Our approach is inspired by the constructive role of foil theories for the understanding of quantum theory [undefu].
 The strategy in defining foil theories is to embed quantum theory in a larger set of theories by relaxing specific assumptions about the structure of the state space, say, or the form of the time evolution.
 We briefly describe typical examples of foil theories and point out that a systematic investigation of generalised state-update rules has, apparently, not yet been carried out.
-Generalised probabilistic theories (GPTs) [barrett_information_2007, hardy_quantum_2001, janotta_generalized_2014], for example, represent a large and versatile collection of foil theories.
+Generalised probabilistic theories (GPTs) [undefe, undefao, undefav], for example, represent a large and versatile collection of foil theories.
 They are defined by their own sets of states, observables, and rules for system composition, all of which may differ from those of standard quantum theory.
 GPTs have become a standard tool to explore and identify distinctive characteristics of quantum theory which they contain as a special case.
-Many GPTs share properties such as the no-cloning theorem, uncertainty relations, or teleportation with quantum theory [barnum_cloning_2006, oppenheim_uncertainty_2010].
-Some of them reproduce the standard quantum-mechanical correlations [wright_gleason-type_2019], while others exhibit superquantum correlations, e.g.
-Boxworld [popescu_quantum_1994].
+Many GPTs share properties such as the no-cloning theorem, uncertainty relations, or teleportation with quantum theory [undefd, undefaam].
+Some of them reproduce the standard quantum-mechanical correlations [undefaaac], while others exhibit superquantum correlations, e.g.
+Boxworld [undefaap].
 Ontological models describe a conceptually different class of foil theories not necessarily based on the familiar Hilbert space structure of quantum theory.
 They aim to explain its predictions in terms of an observer’s incomplete knowledge about the ontic state that provides an objective and complete description of the state of a system.
-In these models, non-classical features of quantum theory such as complementarity of observables, superdense coding, interference phenomena, or uncertainty relations may result from epistemic restrictions [spekkens_evidence_2007, bartlett_reconstruction_2012].
-Other foil theories differ from quantum theory in a single key feature only, such as attempts to modify Born’s rule [galvan_generalization_2008, galley_classification_2017].
-The consequences of associating observables with non-Hermitian operators have been investigated, using PT-symmetric [bender_pt_2024] or normal [roberts_observables_2018] operators, for example.
-Quantum-like theories defined over the real numbers [stueckelberg_quantum_1960, renou_quantum_2021] or quaternions [finkelstein_foundations_1962] rather than the field of complex numbers have also been studied.
-Nonlinear generalisations of Schrödinger’s equation have been suggested [beretta_quantum_1984, abrams_nonlinear_1998, ferrero_nonlinear_2004, kent_nonlinearity_2005, rembielinski_nonlinear_2020], most notably in Weinberg’s work [weinberg_testing_1989].
+In these models, non-classical features of quantum theory such as complementarity of observables, superdense coding, interference phenomena, or uncertainty relations may result from epistemic restrictions [undefaau, undeff].
+Other foil theories differ from quantum theory in a single key feature only, such as attempts to modify Born’s rule [undefal, undefaj], or the tensor product for system composition [undefad].
+The consequences of associating observables with non-Hermitian operators have been investigated, using PT-symmetric [undefj] or normal [undefaas] operators, for example.
+Quantum-like theories defined over the real numbers [undefaax, undefaar] or quaternions [undefaf] rather than the field of complex numbers have also been studied.
+Nonlinear generalisations of Schrödinger’s equation have been suggested [undefm, undefa, undefae, undefay, undefaaq], most notably in Weinberg’s work [undefaaaa].
 Gisin et al.
-[gisin_weinbergs_1990, simon_no-signaling_2001, bassi_no-faster-than-light-signaling_2015] pointed out that they imply violations of the no-signalling principle, although some specific nonlinear transformations were later found not to be ruled out by Gisin’s argument [czachor_nonlocal_1998, kent_nonlinearity_2005, ferrero_nonlinear_2004, helou_extensions_2017, rembielinski_nonlinear_2020].
-In the specific context of modifications to the projection postulate, “causal quantum theory[kent_causal_2005] has been proposed, positing that the Lüders collapse is a well-defined physical process that satisfies strict local causality.
-Additionally, Kent introduced a new type of hypothetical measurement that is capable of revealing complete or partial information about the “local” state of a system, without actually inducing a state update [kent_nonlinearity_2005, kent_quantum_2021, kent_measurements_2023], an idea first presented in [busch_is_1997].
+[undefam, undefaat, undefg] pointed out that they imply violations of the no-signalling principle, although some specific nonlinear transformations were later found not to be ruled out by Gisin’s argument [undefx, undefay, undefae, undefar, undefaaq].
+In the specific context of modifications to the projection postulate, “causal quantum theory[undefaw] has been proposed, positing that the Lüders collapse is a well-defined physical process that satisfies strict local causality.
+Additionally, Kent introduced a new type of hypothetical measurement that is capable of revealing complete or partial information about the “local” state of a system, without actually inducing a state update [undefay, undefaz, undefax], an idea first presented in [undefp].
 Such a measurement would, effectively, allow for yet other nonlinear state transformations without enabling superluminal signalling.
@@
   • ($\mathsf{B}$)
-    The probability that a measurement of $M$ performed on a system in the state $\rho$ yields the outcome represented by $P_{x}$ is given by the Born rule: $\text{prob}(x)=\text{Tr}\left[P_{x}^{A}\,\rho\right]$.
+    The probability that a measurement of $M$ performed on a system in the state $\rho$ yields the outcome represented by $P_{x}$ is given by the Born rule: $\text{prob}(x)=\text{Tr}\left[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}P_{x}}\,\rho\right]$.
 Finally, the projection postulate describes what happens to the state of a quantum system when an observable is measured.
@@
 The trace of the post-measurement state, $\text{Tr}(\omega^{\mathsf{L}}(P_{x},\rho))=\text{Tr}(P_{x}\rho)$, yields the outcome probability as prescribed by the Born rule of postulate ($\mathsf{B}$).
-In addition, the maps (1) and Eq.
-(2) do not require the input state to be normalised.
+In addition, the maps (1) and (2) do not require the input state to be normalised.
 Replacing $\rho$ by $\lambda\rho$, $\lambda\in[0,1]$, shows that the Lüders rule is 1-homogeneous: $\omega^{\mathsf{L}}(P_{x},\lambda\rho)=\lambda\omega^{\mathsf{L}}(P_{x},\rho)$.
 The alternative update rules $\omega(P_{x},\rho)$ considered in Sec.
-5 will be modeled on the Lüders rule in the sense that they are 1-homogeneous, only depend on the measurement outcome $P_{x}$ and the initial state $\rho$, and agree with the Born rule for outcome probabilities.
+5 will be modelled on the Lüders rule in the sense that they are 1-homogeneous, only depend on the measurement outcome $P_{x}$ and the initial state $\rho$, and agree with the Born rule for outcome probabilities.
+[^4] Postulate $(\mathsf{B})$ is understood to assign outcome probabilities to each individual measurement.
+Consequently, Born’s rule on its own does not account for correlations between outcomes of multiple measurements.
+Instead, the update rule is also responsible for the joint probabilities for the outcomes of measurements carried out sequentially, or by distinct parties in composite systems.
+This subtle point proves significant when examining modifications of the Lüders rule $(\mathsf{L})$; see Appendix A for further discussion.
 Further physically motivated properties that any acceptable update rule should satisfy will be discussed in Sec.
@@
 If the outcome $P_{x}^{A}$ is non-degenerate, then the Lüders rule assigns a pure post-measurement state $\omega^{\mathsf{L}}(P_{x}^{A},\rho_{A})$ to subsystem $A$.
-Consequently, the post-measurement state of the composite system must be a product state, whatever the initial state $\rho_{AB}$.
+Consequently, the post-measurement state of the composite system must be a product state, regardless of the initial state $\rho_{AB}$.
 On its own, the projection postulate ($\mathsf{L}$) provides, however, no information about the local state at $B$ upon performing a local measurement of $M_{A}$ on the constituent $A$.
-In principle, any density matrix $\rho_{B}=\text{Tr}_{B}(\rho_{AB})\in\mathcal{S}(\mathcal{H}_{B})$ is compatible.
+In principle, any density matrix $\rho_{B}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\text{Tr}_{A}}(\rho_{AB})\in\mathcal{S}(\mathcal{H}_{B})$ is compatible.
 For example, one could imagine that any local measurement at $A$ throws the state of the composite system into a product state, with the post-measurement state at $B$ being equal to the maximally mixed state $\rho_{B}\propto\mathbb{I}_{B}$.
@@
 Context-independence asserts that the observed outcome and the pre-measurement state alone determine the post-measurement state.
-It introduces a form of non-contextuality [kochen_problem_1967] that is present in quantum theory—namely, that all operationally equivalent experimental runs (i.e.
+It introduces a form of non-contextuality [undefaac] that is present in quantum theory—namely, that all operationally equivalent experimental runs (i.e.
 involving the same initial state and the same measurement outcome) yield identical post-measurement states.
@@
 (A1-A2)
-Completeness &\&& context-independence A complete and context-independent update rule maps any combination of a (local) outcome, $P_{x}^{A}\in\mathcal{P}(\mathcal{H}_{A})$, and a sub-normalised state, $\rho_{AB}\in\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$, to another sub-normalised state,
+Completeness &\& context-independence A complete and context-independent update rule maps any combination of a (local) outcome, $P_{x}^{A}\in\mathcal{P}(\mathcal{H}_{A})$, and a sub-normalised state, $\rho_{AB}\in\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$, to another sub-normalised state,
 $\omega_{AB}:\mathcal{P}(\mathcal{H}_{A})\times\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})\to\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})\,.$ (8)
@@
 (7)); and (iii) assumptions A1-A6 (Eqs.
-(8)-(13)).
+(8)-(13)), e.g.
+$\omega_{AB}:\mathcal{P}(\mathcal{H}_{A})\times\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})\to\bar{\mathcal{S}}(\mathcal{H}_{A}\otimes\mathcal{H}_{B})$ etc.
 Any update rule chosen to replace the projection postulate of quantum theory will give rise to one specific generalised state-update theory.
@@
 5.2 where different extensions of the single-system Lüders rule to multi-partite systems are presented.
-How do the update rules defined here compare with quantum instruments [heinosaari_mathematical_2011], i.e.
+How do the update rules defined here compare with quantum instruments [undefap], i.e.
 collections $\{\mathcal{I}_{x}\}_{x}$ of quantum operations [^10] Quantum operations are linear, completely positive and trace non-increasing maps defined on the space $\mathcal{L}(\mathcal{H})$ of bounded operators acting on $\mathcal{H}$.
@@
 as a collection $\mathcal{G}=\{(p_{1},\rho_{1}),(p_{2},\rho_{2}),\dots,(p_{n},\rho_{n})\}$ where $p_{i}\geq 0$, $\sum_{i}p_{i}=1$ and $\rho_{i}\in\bar{\mathcal{S}}(\mathcal{H})$ represent pure states or improper mixed states.
-The Gemenge $\mathcal{G}$ describes a system that is prepared in state $\rho_{i}$ with probability $p_{i}$ [busch_quantum_1996].
+The Gemenge $\mathcal{G}$ describes a system that is prepared in state $\rho_{i}$ with probability $p_{i}$ [undefr].
 To any Gemenge there corresponds a unique density operator, $\rho_{\mathcal{G}}=\sum_{i=1}^{n}p_{i}\rho_{i}$.
@@
 For instance, the completeness A1 of $\omega$ guarantees that $\rho_{\mathcal{G}_{x}}$ in (14) is a valid state of the system.
-Similar arguments apply to assumptions A3-6, with the exception of context-independence A2, which requires minor adaptation.
+Similar arguments apply to assumptions A3-6, but not to context-independence A2, which requires minor adaptation.
 In fact, the post-measurement state $\rho_{\mathcal{G}_{x}}$ generally does not depend on the initial mixed state $\rho_{\mathcal{G}}$, but rather on the initial Gemenge $\mathcal{G}$.
@@
 Therefore, $\omega^{\mathsf{L}}_{A,x}$ has a unique linear extension to the set of bounded operators $\mathcal{L}(\mathcal{H}_{A})$ (cf.
-[heinosaari_mathematical_2011]).
+[undefap]).
 The Lüders rule (15) for the composite system $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, conditioned on the local measurement outcome $P_{x}^{A}$, is then obtained by setting
@@
 We will now consider measurements that, instead of updating the quantum state, leave it unchanged.
-[^10] The possibility of non-collapsing measurements was listed by von Neumann [vonneumann_mathematische_1932] as early as 1932 as one of three possible reactions of a physical system to a measurement.
+[^13] The possibility of non-collapsing measurements was listed by von Neumann [undefaaz] as early as 1932 as one of three possible reactions of a physical system to a measurement.
 Outcomes still occur probabilistically as dictated by the Born rule, yet, in analogy with classical theory, these measurements do not disturb the pre-measurement state.
@@
 $\omega^{\mathsf{P}}_{A}\left(P_{x}^{A},\rho_{A}\right)=\text{Tr}\left(P_{x}^{A}\rho_{A}\right)\rho_{A}\,.$ (21)
-Non-collapsing measurements, originally introduced in [aaronson_space_2016], cannot be implemented in quantum theory [busch_is_1997].
+Non-collapsing measurements, originally introduced in [undef], cannot be implemented in quantum theory [undefp].
 Indeed, $\omega^{\mathsf{P}}_{A}$ is not convex-linear over $\bar{\mathcal{S}}(\mathcal{H}_{A})$ which implies that, after a passive measurement, one may be able to distinguish between different preparations of the same mixed state.
@@
 $\mathcal{G}=\{(1/2,|0\rangle\langle 0|),(1/2,|1\rangle\langle 1|)\}\quad\mbox{and}\quad\mathcal{G}^{\prime}=\{(1/2,|+\rangle\langle+|),(1/2,|-\rangle\langle-|)\}\,$, (22)
-where |±$|\pm\rangle=(|0\rangle\pm|1\rangle)/\sqrt{2}$.
+where $|\pm\rangle=(|0\rangle\pm|1\rangle)/\sqrt{2}$.
 Both preparations correspond to the maximally mixed state, $\rho_{\mathcal{G}}=\rho_{\mathcal{G}^{\prime}}=\mathbb{I}_{2}/2$.
@@
 Therefore, the non-collapsing update rule $\omega^{\mathsf{P}}_{A}$ does not satisfy preparation indistinguishability, indicating that density operators do not provide complete descriptions of individual systems.
-While complete positivity can be defined for both linear and nonlinear transformations [ando_non-linear_1986], it has been described as “physically unfitting” for nonlinear dynamics [czachor_complete_1998].
+While complete positivity can be defined for both linear and nonlinear transformations [undefb], it has been described as “physically unfitting” for nonlinear dynamics [undefy].
 In fact, although $\omega^{\mathsf{P}}_{A}$ (or, more precisely, extensions of $\omega^{\mathsf{P}}_{A}$ to $\mathcal{L}(\mathcal{H}_{A})$) is not completely positive, the update rule can still be consistently extended to composite systems.
@@
 We called this theory passive quantum theory (pQT).
-An overview of pQT and its key properties has been provided in [fiorentino_quantum_2023].
+An overview of pQT and its key properties has been provided in [undefag].
 The absence of any measurement-induced state update allows, in principle, for tomographic reconstruction of the state of an individual system.
-Thus, the cloning of arbitrary states becomes possible and makes computation in pQT more efficient (in some sense) than in standard quantum theory [busch_is_1997, kent_nonlinearity_2005].
+Thus, the cloning of arbitrary states becomes possible and makes computation in pQT more efficient (in some sense) than in standard quantum theory [undefp, undefay].
 The passive update rule of pQT also satisfies ideality but is not locally tomographic and does not allow for violations of Bell’s inequalities (cf.
@@
 As a result, correlations between measurements on different subsystems cannot be unambiguously determined.
-The projection postulate introduced by von Neumann [vonneumann_mathematische_1932] is a notable example of state-update that does not define a valid update rule.
+The projection postulate introduced by von Neumann [undefaaz] is a notable example of state-update that does not define a valid update rule.
 For non-degenerate measurements, von Neumann’s and Lüders’ postulates coincide.
@@
 [^16] One could, however, modify (36) into a valid update rule by, for example, fixing a preferred set $\mathcal{B}_{x}$ for each experiment yielding the outcome $P^{A}_{x}$, and subsequently extend the map to composite systems via complete positivity.
-Finally, we present a simple example of an update rule that violates local covariance A4.
-Assume that any measurement on a system causes the pre-measurement state to undergo a fixed, observer-independent and non-trivial [^14] If $U=\mathbb{I}$, the passive measurements of Sec.
-5.3 result consistent with A4.
+Finally, we present a simple example of an update rule that violates local covariance A3.
+Assume that any measurement on a system causes the pre-measurement state to undergo a fixed, observer-independent and non-trivial [^17] If $U=\mathbb{I}$, we recover the passive measurements of Sec.
+5.3, which are consistent with A3.
 unitary transformation $U\in\mathcal{U}(\mathcal{H})$,
@@
 5.3) provides another example of an update rule satisfying ideality.
-Lüders presents two main objections to von Neumann’s projection postulate [luders_uber_1950].
+Lüders presents two main objections to von Neumann’s projection postulate [undefaaf].
 Firstly, he argues that the post-measurement state should depend only on the outcome and the initial state.
@@
 “the measurement of a highly degenerate quantity permits only relatively weak assertions regarding the considered ensemble.
-For that reason, the resulting change in state should likewise be small” ([luders_concerning_2006], p.
+For that reason, the resulting change in state should likewise be small” ([undefaae], p.
 665).
@@
 In the limiting case of $P^{A}_{y}=\mathbb{I}_{A}$, one obtains $\omega_{A}\left(\mathbb{I}_{A},\rho_{A}\right)\propto\rho_{A}$, in agreement with Lüders’ remark.
-In some form or another, the assumption of coherence has appeared already in the literature, both in the context of quantum theory [bell_moral_1966, herbut_compatibility_2004, khrennikov_neumann_2009] and of generalised probabilistic theories [kleinmann_sequences_2014].
-For measurements on non-composite quantum systems, coherence entails both deterministic repeatability [kleinmann_sequences_2014] and ideality.
+In some form or another, the assumption of coherence has appeared already in the literature, both in the context of quantum theory [undefi, undefat, undefaaa] and of generalised probabilistic theories [undefaab].
+For measurements on non-composite quantum systems, coherence entails both deterministic repeatability [undefaab] and ideality.
 Lemma 1.
@@
 the original projection postulate ($\mathsf{L}$).
-This argument was first presented in 1966 [bell_moral_1966] and similar versions of it have since been rediscovered several times; see Appendix A for a summary.
+This argument was first presented in 1966 [undefi] and similar versions of it have since been rediscovered several times; see Appendix A for a summary.
 Below, we present a simple proof of this result.
@@
 3)—see Theorem 2 below.
-This final step in deriving the projection postulate is often overlooked [bell_moral_1966, herbut_compatibility_2004, martinez_search_1990], as the extension of $\omega^{\mathsf{L}}_{A}$ to local measurements on composite systems is typically assumed implicitly, despite not being stated in or implied by the standard postulates.
+This final step in deriving the projection postulate is often overlooked [undefi, undefat, undefaai], as the extension of $\omega^{\mathsf{L}}_{A}$ to local measurements on composite systems is typically assumed implicitly, despite not being stated in or implied by the standard postulates.
 Theorem 2.
@@
 Any valid state-update rule must satisfy six core requirements which include the uniqueness of post-measurement states, no-signalling, and consistency for measurements carried out by observers with access to subsystems only.
-We have shown that quantum theory and passive quantum theory [fiorentino_quantum_2023] satisfy the requirements.
+We have shown that quantum theory and passive quantum theory [undefag] satisfy the requirements.
 Three new examples of consistent generalised state-update rules have been defined, each one leading to a toy theory structurally different from quantum theory.
@@
 This result agrees with earlier ones based on similar concepts describing a form of information-disturbance trade-off.
-However, our systematic approach to GURs highlights the fact that known derivations of the Lüders rule are incomplete since they typically do not take composite systems into account.
+However, our systematic approach to GURs highlights the fact that many derivations of the Lüders rule may be considered incomplete by not explicitly addressing composite systems.
 We find that coherence on its own is insufficient to single out the projection postulate when considering composite systems; equally, composition compatibility alone is not strong enough to do so.
@@
 Importantly, complete positivity is not strong enough to ensure that such measurement-induced transformations assign post-measurement states to composite systems consistently.
-In the words of [czachor_complete_1998], complete positivity is “physically unfitting” for this purpose.
-Reconstructions of quantum theory that engage—either directly or indirectly—with post-measurement states usually assume measurement-induced state transformations to be linear in the pre-measurement state [barrett_information_2007, kleinmann_sequences_2014, masanes_measurement_2019].
-This assumption, often left implicit [stacey_masanes-galley-muller_2022, kent_measurements_2023], excludes a priori certain types of measurement behaviours—such as the trivial non-collapsing rule of “passive quantum theory”—that are, in fact, operationally well-defined and consistent with fundamental principles such as no-signalling (see Sec.
+In the words of [undefy], complete positivity is “physically unfitting” for this purpose.
+Reconstructions of quantum theory that engage—either directly or indirectly—with post-measurement states usually assume measurement-induced state transformations to be linear in the pre-measurement state [undefe, undefaab, undefaak].
+This assumption, often left implicit [undefaaw, undefax], a priori excludes measurement behaviours such as the trivial non-collapsing rule of “passive quantum theory” that are, in fact, operationally well-defined and consistent with fundamental principles such as no-signalling (see Sec.
 4.2).
-Preliminary results based on the framework of GURs (as well as earlier works, e.g.
-[flatt_gleason-busch_2017]) suggest that the linearity of update rules could be derived from other, more natural operational assumptions.
+Preliminary results based on the framework of GURs and earlier work (e.g.
+[undefaan, undefah]) suggest that the linearity of update rules could be derived from natural operational assumptions, mirroring efforts to establish the linearity of the quantum time evolution [undefam, undefaaab].
 It seems promising to investigate the role of nonlocality from the perspective of theories with generalised update rules.
@@
 Acknowledgements
-The authors are grateful for support through the Leverhulme grant RPG-2024-201.
-Appendix A Summary of derivations of the Lüders rule
+The authors are grateful for support through grant RPG-2024-201 by the Leverhulme Trust.
+Appendix A Derivations of the Lüders rule – A brief survey
 Derivations of the Lüders rule have a long history, with similar arguments being made over the years.
@@
 A key observation is that most of the justifications focus on non-composite systems, either ignoring measurements on composite systems or making implicit assumptions about them.
-In 1966, Bell and Nauenberg showed that the quantum mechanical state-update can be obtained by assuming that the outcome probability distribution of a common refinement for commuting observables $M_{1}$ and $M_{2}$ equals the joint probability distribution obtained by measuring $M_{1}$ and $M_{2}$ sequentially, in any order [bell_moral_1966].
+In 1966, Bell and Nauenberg showed that the quantum mechanical state-update can be obtained by assuming that the outcome probability distribution of a common refinement for commuting observables $M_{1}$ and $M_{2}$ equals the joint probability distribution obtained by measuring $M_{1}$ and $M_{2}$ sequentially, in any order [undefi].
 This assumption is mathematically equivalent to that of coherence (Def.
@@
 6.
-To the best of our knowledge, the only reference to this early operational derivation of the Lüders rule in non-composite systems is by Herbut [herbut_compatibility_2004].
+To the best of our knowledge, the only reference to this early operational derivation of the Lüders rule in non-composite systems is by Herbut [undefat].
 Over the last 40 years, the argument by Bell and Nauenberg has been rediscovered at least three times, using the concept of coherence.
@@
 To do so, they describe the state-update following measurements in terms of a generalised conditional probability defined on the generalised probability space—i.e.
-a probability measure on $\mathcal{P}(\mathcal{H})$ compatible with Born’s rule and satisfying “coherence” [cassinelli_conditional_1983].
-In 2009, Khrennikov showed that the Lüders rule for single systems follows from von Neumann’s projection postulate for non-degenerate observables if “coherence” is assumed [khrennikov_neumann_2009].
+a probability measure on $\mathcal{P}(\mathcal{H})$ compatible with Born’s rule and satisfying “coherence” [undeft].
+In 2009, Khrennikov showed that the Lüders rule for single systems follows from von Neumann’s projection postulate for non-degenerate observables if “coherence” is assumed [undefaaa].
 Claiming that “this important observation remained unnoticed for the past 70 years”, the author appears to be unaware of Bell and Nauenberg’s work.
-In 2014, Kleinmann developed a formalism for sequential measurements in a broad class of generalised probabilistic theories (GPTs), introducing a generalisation of the Lüders rule based on coherence [kleinmann_sequences_2014].
+In 2014, Kleinmann developed a formalism for sequential measurements in a broad class of generalised probabilistic theories (GPTs), introducing a generalisation of the Lüders rule based on coherence [undefaab].
 A simple derivation of the non-composite Lüders rule based on coherence is also given, without reference to earlier work.
@@
 5.1), thereby excluding operationally valid alternatives such as the passive update rule from the outset.
-Herbut actually presents two distinct derivations applying to measurements on non-composite systems in a paper from 1969 [herbut_derivation_1969].
+Herbut actually presents two distinct derivations applying to measurements on non-composite systems in a paper from 1969 [undefas].
 In an information-theoretic approach, he postulates (i) deterministic repeatability (see Sec.
@@
 The Lüders rule is then shown to map the pre-measurement state to the least distinguishable state compatible with outcome repeatability.
-This result remained largely unnoticed, while anticipating later arguments deriving the state-update from minimisation principles, including those by Marchand [marchand_statistical_1977], Dieks [dieks_distance_1983], and Hadjisavvas [hadjisavvas_distance_1981].
-More recently, the single-system Lüders rule was obtained from minimising non-symmetric quantum information distances [kostecki_luders_2014, hellmann_quantum_2016], thereby establishing a mathematical connection to the Bayes-Laplace rule.
-Ozawa [ozawa_quantum_1984, ozawa_quantum_1998] derives the update rule from Bayes’ rule by assuming the standard joint probability formula for local quantum measurements.
-Herbut’s second derivation in [herbut_derivation_1969] takes an operational approach.
+This result remained largely unnoticed, while anticipating later arguments deriving the state-update from minimisation principles, including those by Marchand [undefaag], Dieks [undefab], and Hadjisavvas [undefan].
+More recently, the single-system Lüders rule was obtained from minimising non-symmetric quantum information distances [undefaad, undefaq], thereby establishing a mathematical connection to the Bayes-Laplace rule.
+Ozawa [undefaan, undefaao] derives the quantum instrument formalism describing state-updates from Bayes’ rule by assuming the standard joint probability formula for local quantum measurements.
+In a sense, the author adopts a stronger interpretation of postulate $(\mathsf{B})$ (cf.
+Sec.
+3.1) than the one used explicitly by us, and implicitly by other authors [undefay, undef, undefaz, undefax] where $(\mathsf{L})$ is modified.
+The stronger interpretation of Born’s rule is assumed to govern individual measurement outcomes as well as correlations between local measurements carried out by different experimenters with access to parts of a composite system only.
+Essentially, it subsumes the property of local tomography (cf.
+Sec.
+5.1).
+Herbut’s second derivation in [undefas] takes an operational approach.
 In addition to deterministic repeatability, it is assumed that performing an unconditional measurement of observable $M$ does not affect the expectation values of any observables commuting with $M$.
-The argument focuses on measurements on single systems.
+The argument focusses on measurements on single systems.
 However, Herbut also briefly considers composite systems, discussing in particular how local Lüders measurements modify joint states without enabling signalling.
@@
 What is more, Herbut takes convex-linearity over states to be a fundamental requirement for state updates, without justifying this assumption.
-Herbut revisits the question in a 1974 paper [herbut_minimal_1974].
+Herbut revisits the question in a 1974 paper [undefau].
 Here, the Lüders rule for single systems is obtained by combining the assumption of deterministic repeatability with the “preservation of sharp values”.
@@
 When proving this result for improper mixtures, Herbut implicitly adopts “composition compatibility” to describe local measurements.
-Martinez [martinez_search_1990] offers an early account of known derivations of the Lüders rule including Herbut’s work but not Bell and Nauenberg’s contribution.
-The paper also demonstrates that the converse of Lüders theorem [luders_uber_1950, busch_quantum_1996] implies the quantum mechanical collapse for non-composite systems: if, for any pair of commuting observables $M_{1}$ and $M_{2}$ and any initial state, we assume that the sequence of measurements $M_{1}\prec M_{2}\prec M_{1}$ always produces identical outcomes for $M_{1}$ regardless of the outcome of the $M_{2}$ measurement, then the post-measurement state must necessarily be given by the Lüders rule.
-Martinez further examines a quantum-logical derivation employing the “Sasaki hook” presented in [friedman_quantum_1978].
+Martinez [undefaai] offers an early account of known derivations of the Lüders rule including Herbut’s work but not Bell and Nauenberg’s contribution.
+The paper also demonstrates that the converse of Lüders theorem [undefaaf, undefr] implies the quantum mechanical collapse for non-composite systems: if, for any pair of commuting observables $M_{1}$ and $M_{2}$ and any initial state, we assume that the sequence of measurements $M_{1}\prec M_{2}\prec M_{1}$ always produces identical outcomes for $M_{1}$ regardless of the outcome of the $M_{2}$ measurement, then the post-measurement state must necessarily be given by the Lüders rule.
+Martinez further examines a quantum-logical derivation employing the “Sasaki hook” presented in [undefai].
 The extension of the update rule to local measurements on subsystems is not considered.
-The proof that establishes the Lüders instrument as the unique ideal quantum instrument for sharp discrete measurements dates back to Davies [davies_quantum_1976], as recognised in [busch_quantum_1996, busch_repeatable_1995].
+The proof that establishes the Lüders instrument as the unique ideal quantum instrument for sharp discrete measurements dates back to Davies [undefz], as recognised in [undefr, undefq].
 The notion of a quantum instrument incorporates two key assumptions: convex-linearity of the single-system state-update rule, and the standard linear extension to composite systems enabled by the complete positivity of $\omega_{A}$ (see Eq.
@@
 5.1).
-This argument for isolating the Lüders collapse for single systems has been rediscovered occasionally (see  [marinkovic_note_1983], for example).
+This argument for isolating the Lüders collapse for single systems has been rediscovered occasionally (see [undefaah], for example).
 In 2019, Masanes et al.
-advance the argument that both the Born rule and the post-measurement state rule can be derived from the other postulates of quantum theory when supplemented by suitable operational constraints [masanes_measurement_2019].
+advance the argument that both the Born rule and the post-measurement state rule can be derived from the other postulates of quantum theory when supplemented by suitable operational constraints [undefaak].
 In doing so, they consider alternative rules for computing outcome probabilities, explicitly discussing both composite systems and local measurements.
 Having established the uniqueness of the Born rule within their framework, they assert that quantum instruments constitute the only consistent description of state updates.
-This conclusion, along with their broader framework, has attracted some criticism [stacey_masanes-galley-muller_2022, galley_reply_2022, kent_measurements_2023, masanes_response_2023].
-The argument for quantum instruments depends on an assumption regarding sequential measurements that, as already shown in [flatt_foundations_2019], is mathematically equivalent to requiring convex-linearity of state updates.
-Conceptually, Masanes et al.’s approach mirrors Kleinmann’s framework [kleinmann_sequences_2014] in that it is limited a priori to update rules satisfying preparation indistinguishability.
+This conclusion, along with their broader framework, has attracted some criticism [undefaaw, undefak, undefax, undefaaj, undefaav].
+The argument for quantum instruments depends on an assumption regarding sequential measurements that, as already shown in [undefah], is mathematically equivalent to requiring convex-linearity of state updates.
+Conceptually, Masanes et al.’s approach mirrors Kleinmann’s framework [undefaab] in that it is limited a priori to update rules satisfying preparation indistinguishability.
 This restriction explains why they adopt the completely positive extension (18) for local measurements on composite systems: any other extension would necessarily violate their assumption of preparation indistinguishability.
@@
 Theorem 2) are required to exclude alternative update rules in the multi-partite case.
-None of these properties are contained in the postulates ($\mathsf{S}$) to ($\mathsf{B}$) of quantum theory.
-[…]
-[heading=bibintoc]